### v1

Abstract
In this paper we introduce the $q$-differential operator for $q$-spinor variables. We establish the $q$-spinor chain rule , the new $q$-differential operator, the $q$-Dirac differential operators and the $q$-complex spinor integrals. We also define the $q$-spinor differential equation. The suggestions for further work at the end of the paper.
Keywords:
$q$-Differential operators, $q$-Dirac operator, $q$-spinor variables, integral formulas in $q$-spinor variables, differential equation in $q$-spinor variables.
msclass: 81Q99, 46E99, 35A24, 15A66 , 16T99 . 17B37.
1 Introduction
Based on the work of Beretetskii et al., Lachieze-Rey, Gori et al., and Cartan, the spinor $\psi^{\alpha}$ is defined as a magnitude components $\alpha=1,2$ expressed as $\psi^{\alpha}=\begin{bmatrix}\psi_{1}\\ \psi_{2}\end{bmatrix}$ and its complex conjugate $\dot{\varphi}$ in terms of the rotation matrices (see [2],[1],[9],[5] for more details). Based on the work of the previously mentioned authors, there are two types of operations on spinors, which are reflections and rotations. In group theory, the set of rotations described by the matrices with complex entries is group $SU(2)$, whose generators are the Pauli matrices, described in the work of Zettili [17]. With respect to rotation matrices, Gori et al. mention, in their work, the rotation matrices that originated the Pauli matrices in the form:
$R_{x}(\theta)=\begin{bmatrix}\cos(\theta/2)&i\sin(\theta/2)\\ i\sin(\theta/2)&\cos(\theta/2)\end{bmatrix},R_{y}(\theta)=\begin{bmatrix}\cos(\theta/2)&\sin(\theta/2)\\ \sin(\theta/2)&\cos(\theta/2)\end{bmatrix},R_{z}(\theta)=\begin{bmatrix}e^{i\theta/2}&0\\ 0&e^{-i\theta/2}\end{bmatrix},$ (1)
being $\theta$ the angle of rotation [5]. Beretetskii et al. define the covariance and contravariance over the spinors by the relation $\psi^{\prime 1}=\psi_{2},\psi^{\prime 2}=-\psi_{1}$ from the matrix $g_{\alpha\beta}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ and, similarly, for pointed spinors, $\psi_{\dot{1}}=\psi^{\dot{2}},\psi_{\dot{2}}=-\psi^{\dot{1}}$ [1]. The same author defines bispinors as the pair $(\psi^{\alpha},\varphi_{\dot{\alpha}})$, which form a broader group of Lorentz, and, with them, the scalar product is formulated as $(\psi^{\alpha},\varphi_{\dot{\alpha}})\cdot(f^{\alpha},h_{\dot{\alpha}})$. The author of the reference [9] mentions algebra $Cl(3)$ as a space-time formulation generated by vectors $e_{\mu}$, which form a basis for $\mathds{R}^{1,3}$ that satisfies the relation $e_{\mu}\cdot e_{\nu}=g_{\mu\nu}$, inducing a 16 -membered basis, as described below:
  • 1
    $1$ scalar,
  • 2
    $(e_{0},e_{1},e_{2},e_{3})$ 4-vector,
  • 3
    $(e_{0}e_{1},e_{0}e_{2},e_{0}e_{3},e_{1}e_{2},e_{2}e_{3},e_{3}e_{1})$ 6- bivectors,
  • 4
    $(e_{1}e_{2}e_{3},e_{0}e_{2}e_{3},e_{0}e_{1}e_{3},e_{0}e_{1}e_{2})$ 4-trivectors,
  • 5
    $e_{5}\equiv e_{0}e_{1}e_{2}e_{3}$ pseudoescalar.
In accordance with the above, the same author describes the Weyl spinors as
$\begin{bmatrix}\psi^{1}\\ \psi^{2}\end{bmatrix}$ and the Dirac spinor $\begin{bmatrix}\psi^{\alpha}\\ \varphi^{\dot{\alpha}}\end{bmatrix}$, as defined in the work of Beretetskii et al [1].
The $q$ Lorentzian algebra was defined in the reference [12]. The quantum complex spinors have components $\psi_{1}$ and $\psi^{2}$ and conjugates $\varphi^{\dot{1}}$ and $\varphi_{\dot{2}}$. For all $q\in\mathds{R}-\left\{0\right\}$, they satisfy the following $q$ relations
$\psi_{1}\psi^{2}=q\psi^{2}\psi_{1},$ $\ \psi^{2}\varphi^{\dot{1}}=q\varphi^{\dot{1}}\psi^{2},$ (2)
$\psi_{1}\varphi^{\dot{1}}=\varphi^{\dot{1}}\psi_{1}-q(q+q^{-1})^{1/2}\varphi_{\dot{2}}\varphi^{\dot{1}},$ $\ \psi^{2}\varphi^{\dot{1}}=\varphi^{\dot{1}}\psi^{2},$ (3)
$\psi_{1}\varphi_{\dot{2}}=q\varphi_{\dot{2}}\psi_{1},$ $\ \varphi^{\dot{1}}\varphi_{\dot{2}}=q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}}.$ (4)
Considering spinors $\psi^{\alpha},\varphi^{\dot{\alpha}}$ and $\left\{\tau^{1},T^{2},S^{1},\sigma^{2}\right\}$ as the generators of the $q$ Lorentzian algebra for the group $U_{q}(su(2))$ [13], we have:
  • 1.
    For $\psi^{\alpha},\alpha=1,2$
    $\tau^{1}\psi_{1}=$ $\ \psi_{1}\tau^{1},$ (5)
    $\tau^{1}\psi^{2}=$ $\ \psi^{2}\tau^{1}-q(q+q^{-1})^{2}\psi_{1}T^{2},$ (6)
    $T^{2}\psi_{1}=$ $\ q^{-1}\psi_{1}T^{2},$ (7)
    $S^{1}\psi_{1}=$ $\ q\psi_{1}S^{1},$ (8)
    $T^{2}\psi^{2}=$ $\ q\psi^{2}T^{2},$ (9)
    $S^{1}\psi^{2}=$ $\ q^{-1}\psi^{2}S^{1}-\psi_{1}\sigma^{2},$ (10)
    $\sigma^{2}\psi_{1}=$ $\ \psi_{1}\sigma^{2},$ (11)
    $\sigma^{2}\psi^{2}=$ $\ \psi^{2}\sigma^{2},$ (12)
  • 2.
    Their complex conjugates $\varphi^{\dot{\alpha}},\alpha=\dot{1},\dot{2}$
    $\tau^{1}\varphi^{\dot{1}}=$ $\ q^{-1}\varphi^{\dot{1}}\tau^{1},$ (13)
    $\tau^{1}\varphi_{\dot{2}}=$ $\ q\varphi_{\dot{2}}\tau^{1},$ (14)
    $T^{2}\varphi^{\dot{1}}=$ $\ \varphi^{\dot{1}}T^{2}+q^{-1}\varphi_{\dot{2}}\tau^{1},$ (15)
    $T^{2}\varphi_{\dot{2}}=$ $\ \varphi_{\dot{2}}T^{2},$ (16)
    $S^{1}\varphi^{\dot{1}}=$ $\ \varphi^{\dot{1}}S^{1},$ (17)
    $\sigma^{2}\varphi^{\dot{1}}=$ $\ q\varphi^{\dot{1}}\sigma^{2}+(q+q^{-1})^{2}\varphi_{\dot{2}}S^{1},$ (18)
    $\sigma^{2}\varphi_{\dot{2}}=$ $\ q\varphi_{\dot{2}}\sigma^{2}.$ (19)
Deformed commutation relations for $q$- Lorentzian algebra are defined in the next proposition, on the quantum-symmetric plane and the quantum anti-symmetric plane.
Proposition 1.1.
Consider generator $T$ of the set $\left\{\tau^{1},T^{2},S^{1},\sigma^{2}\right\}$ for the algebra $U_{q}(su(2))$ and the relations 2, 3, and 4 , defined in [12] and [13] . The $q$ Lorentzian algebra for spinors in the deformed space is defined through the following relations:
$T(\psi_{1}\psi^{2}-q\psi^{2}\psi_{1})=$ $\ (\psi_{1}\psi^{2}-q\psi^{2}\psi_{1})T,$ (20)
$T(\psi_{1}\psi^{2}-q\psi^{2}\psi_{1})=$ $\ (\psi_{1}\psi^{2}-q\psi{2}\psi^{1})T,$ (21)
$T(\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})=$ $\ (\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})T,$ (22)
$T(\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})=$ $\ (\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})T.$ (23)
Definition 1.2.
The following are the bosonic $q$ deformed Minkowskian Pauli spin matrices defined in the Schmidt work [13]:
$(\sigma^{+})_{\alpha\dot{\beta}}=\begin{bmatrix}0&0\\ 0&q\end{bmatrix},\quad(\sigma^{-})_{\alpha\dot{\beta}}=\begin{bmatrix}q&0\\ 0&0\end{bmatrix},$ (24)
$(\sigma^{3})_{\alpha\dot{\beta}}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\sigma^{0})_{\alpha\dot{\beta}}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&-q^{-1/2}\\ q^{1/2}&0\end{bmatrix}.$ (25)
Likewise, the conjugated Pauli matrices are:
$(\overline{\sigma}^{+})_{\dot{\alpha}\beta}=\begin{bmatrix}0&0\\ 0&q^{-1}\end{bmatrix},\quad(\overline{\sigma}^{-})_{\dot{\alpha}\beta}=\begin{bmatrix}q^{-1}&0\\ 0&0\end{bmatrix},$ (26)
$(\overline{\sigma}^{3})_{\dot{\alpha}\beta}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\overline{\sigma}^{0})_{\dot{\alpha}\beta}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{-1/2}\\ -q^{1/2}&0\end{bmatrix}.$ (27)
The inverse Pauli matrices
$(\sigma^{-1}_{+})_{\alpha\dot{\beta}}=\begin{bmatrix}0&0\\ 0&q^{-1}\end{bmatrix},\quad(\sigma^{-1}_{-})_{\alpha\dot{\beta}}=\begin{bmatrix}q^{-1}&0\\ 0&0\end{bmatrix},$ (28)
$(\sigma^{-1}_{3})_{\alpha\dot{\beta}}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\sigma^{-1}_{0})_{\alpha\dot{\beta}}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&-q^{-1/2}\\ q^{1/2}&0\end{bmatrix}.$ (29)
Finally
$(\overline{\sigma}^{-1}_{+})_{\dot{\alpha}\beta}=\begin{bmatrix}0&0\\ 0&q^{-1}\end{bmatrix},\quad(\overline{\sigma}^{-1}_{-})_{\dot{\alpha}\beta}=\begin{bmatrix}q^{-1}&0\\ 0&0\end{bmatrix},$ (30)
$(\overline{\sigma}^{-1}_{3})_{\dot{\alpha}\beta}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\overline{\sigma}^{-1}_{0})_{\dot{\alpha}\beta}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&-q^{-1/2}\\ q^{1/2}&0\end{bmatrix}.$ (31)
Definition 1.3.
The $q$ Lorentzian spinor variables or $q$ spinor variables are defined according to the expressions (2), (3), and (4) as follows:
$u^{2}_{1}\equiv$ $\ \psi_{1}\psi^{2}-q\psi^{2}\psi_{1},$ (32)
$v^{\dot{1}2}\equiv$ $\ \psi^{2}\varphi^{\dot{1}}-q\varphi^{\dot{1}}\psi^{2},$ (33)
$x^{\dot{1}}_{1\dot{2}}\equiv$ $\ \psi_{1}\varphi^{\dot{1}}-\varphi^{\dot{1}}\psi_{1}+q(q+1)^{-1/2}\varphi_{\dot{2}}\varphi^{\dot{1}},$ (34)
$y^{2\dot{1}}\equiv$ $\ \psi^{2}\varphi^{\dot{1}}-\varphi^{\dot{1}}\psi^{2},$ (35)
$z^{\dot{1}}_{\dot{2}}\equiv$ $\ \psi_{1}\varphi_{\dot{2}}-q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}},$ (36)
$t_{1\dot{2}}\equiv$ $\ \psi_{1}\varphi_{\dot{2}}-q\varphi_{\dot{2}}\psi_{1}.$ (37)
Definition 1.4.
We consider the set $U=\left\{u^{2}_{1},v^{\dot{1}2},x^{\dot{1}}_{1\dot{2}},z^{\dot{1}}_{\dot{2}},y^{2\dot{1}},t_{1\dot{2}}\right\}\subset\mathds{C}$. A function on the $q$ spinor variables is defined as $\Psi(U)=\Psi(u^{2}_{1},v^{\dot{1}2},x^{\dot{1}}_{1\dot{2}},z^{\dot{1}}_{\dot{2}},y^{2\dot{1}},t_{1\dot{2}})$.
Definition 1.5.
Let $f,g:U\longrightarrow\mathds{C}$ be functions and $u^{\beta}\in U$. The following properties are satisfy on the $q$ spinor variables, we state some clear properties of the functions on the $q$ spinor variables
  • 1.
    $(f+g)(u^{\beta})=f(u^{\beta})+g(u^{\beta})$.
  • 2.
    $(f\cdot g)(u^{\beta})=f(u^{\beta})\cdot g(u^{\beta})$.
  • 3.
    $(f-g)(u^{\beta})=f(u^{\beta})-g(u^{\beta})$.
  • 4.
    $\left(\dfrac{f}{g}\right)(u^{\beta})=\dfrac{f(u^{\beta})}{g(u^{\beta})},\quad g(u^{\beta})\neq 0$.
Definition 1.6.
For a function $f:U\longrightarrow\mathds{C}$ and $u^{\beta}\in\mathbb{C}$ , the $q$ spinor derivative is defined as [8] :
$\dfrac{\mathrm{d}_{q}f}{\mathrm{d}_{q}u^{\beta}}=\dfrac{f((qu)^{\beta})-qf(u^{\beta})}{(qu)^{\beta}-qu^{\beta}},$ (38)
and its conjugate complex
$\dfrac{\mathrm{d}_{q}f}{\mathrm{d}_{q}v^{\dot{\alpha}}}=\dfrac{f((qv)^{\dot{\alpha}})-qf(v^{\dot{\alpha}})}{(qv)^{\dot{\alpha}}-qv^{\dot{\alpha}}}.$ (39)
1.1 Clifford algebra and Dirac operator
Let $\left\{\gamma_{1},\gamma_{2},\cdots,\gamma_{n},\right\}$ be an orthonormal basis of $\mathds{R}^{n}$. The Clifford algebra is generated over $\mathbb{R}^{n}$ under the relation
$\gamma_{\mu}\gamma_{\nu}+\gamma_{\nu}\gamma_{\mu}=-2\delta_{\mu\nu}\gamma_{0},\quad\gamma_{\mu}^{2}=-|\gamma_{\mu}|^{2}\gamma_{0},\quad\mu,\nu=1,2,...,n,$ (40)
where $\delta_{\mu\nu}$ is the Kronecker symbol (see [11], [3], [10] for more details). We will denote the Clifford algebra by $Cl_{n}$, and each element in $Cl_{n}$ can be expressed by its components as $\sum_{a}\gamma_{a}x_{a}$, where $a=(\mu_{1},...,\mu_{n})$ with each $\mu_{l}\in\left\{1,2,...,n\right\}$. Any element $\textit{{x}}\in\mathds{R}^{n}$ can be identified with a 1-vector in the Clifford algebra [10]
$(x_{1},x_{2},...,x_{n})\longrightarrow\textit{{x}}=x_{1}\gamma_{1}+x_{2}\gamma_{2}+\cdots+x_{n}\gamma_{n}.$ (41)
On other hand, the Dirac operator used here is
$D:=\sum\limits_{\mu=1}^{n}\gamma_{\mu}\frac{\partial}{\partial x_{\mu}},$ (42)
we refer to reader to [3], [6], [4] for more details.
1.2 $q$-Deformed Dirac matrices
Definition 1.7.
The $q$-deformed Dirac matrices are defined in [15], and are given by
$\gamma_{\mu}:=\ \begin{bmatrix}0&(\sigma_{\mu})^{\alpha}_{\dot{\beta}}\\ (\overline{\sigma}_{\mu})^{\dot{\alpha}}_{\beta}&0\end{bmatrix},$ (43)
where $(\sigma_{\mu})^{\alpha}_{\dot{\beta}}$ and $(\overline{\sigma}_{\mu})^{\dot{\alpha}}_{\beta}$ denote the Pauli matrices of $q-$ deformed Minkowski space (e.g. [14] for more details), and are defined as
$\begin{split}(\sigma_{+})^{\alpha}_{\dot{\beta}}=\begin{bmatrix}0&0\\ 0&kq^{1/2}\lambda^{1/2}_{+}\end{bmatrix},\quad(\sigma_{3})^{\alpha}_{\dot{\beta}}=k\begin{bmatrix}0&q\\ 1&0\end{bmatrix},\\ (\sigma_{-})^{\alpha}_{\dot{\beta}}=k\begin{bmatrix}q^{1/2}\lambda^{1/2}_{+}&0\\ 0&0\end{bmatrix},\quad(\sigma_{0})^{\alpha}_{\dot{\beta}}=k\begin{bmatrix}0&-q^{-1}\\ 1&0\end{bmatrix},\end{split}$ (44)
and their conjugated counterparts
$\begin{split}(\overline{\sigma}_{+})^{\alpha}_{\dot{\beta}}=\begin{bmatrix}0&0\\ 0&\overline{k}q^{-1/2}\lambda^{1/2}_{+}\end{bmatrix},\quad(\overline{\sigma}_{3})^{\alpha}_{\dot{\beta}}=\overline{k}\begin{bmatrix}0&1\\ q^{-1}&0\end{bmatrix},\\ (\overline{\sigma}_{-})^{\alpha}_{\dot{\beta}}=\overline{k}\begin{bmatrix}q^{-1/2}\lambda^{1/2}_{+}&0\\ 0&0\end{bmatrix},\quad(\overline{\sigma}_{0})^{\alpha}_{\dot{\beta}}=\overline{k}\begin{bmatrix}0&1\\ -q&0\end{bmatrix},\end{split}$ (45)
where $k,\overline{k}$ are characteristic parameters associated to bosons ($q=+1$) and fermions ($q=-1$).
1.3 $q$-Spinor complex integral formulas
Definition 1.8.
[8] Let $\Gamma_{q}$ be the closed contour of the deformed quantum complex plane, and $u_{0}^{\beta},v^{\dot{\alpha}}\subset\Gamma_{q}$ point spinors contained in the contour. The $q$- spinor complex integral formulas are defined by
$\oint_{\Gamma_{q}}\dfrac{\Psi(u^{\beta})\mathrm{d}_{q}u^{\beta}}{(qu)^{\beta}-qu^{\beta}_{0}}$ $=\ \dfrac{1}{q}\sum_{n=0}^{\infty}\left[(\overline{\sigma}_{\mu})_{\dot{\alpha}\beta}\Psi(u_{0}^{\beta})\right]^{n},$ (46)
$\oint_{\Gamma_{q}}\dfrac{\Psi((qu)^{\beta})\mathrm{d}_{q}u^{\beta}}{(qu)^{\beta}-qu^{\beta}_{0}}$ $=\ \sum_{n=0}^{\infty}\left[(\overline{\sigma}_{\mu})_{\dot{\alpha}\beta}\Psi((qu_{0})^{\beta})\right]^{n},$ (47)
$\oint_{\Gamma_{q}}\dfrac{\Psi(v^{\dot{\alpha}})\mathrm{d}_{q}v^{\dot{\alpha}}}{(qv)^{\dot{\alpha}}-qv^{\dot{\alpha}}_{0}}$ $=\ \dfrac{1}{q}\sum_{m=0}^{\infty}\left[(\sigma_{\mu})_{\alpha\dot{\beta}}\Psi(v_{0}^{\dot{\alpha}})\right]^{m},$ (48)
$\oint_{\Gamma_{q}}\dfrac{\Psi((qv)^{\dot{\alpha}})\mathrm{d}_{q}v^{\dot{\alpha}}}{(qv)^{\dot{\alpha}}-qv^{\dot{\alpha}}_{0}}$ $=\ \sum_{m=0}^{\infty}\left[(\sigma_{\mu})_{\alpha\dot{\beta}}\Psi((qv_{0})^{\dot{\alpha}})\right]^{m}.$ (49)
Motivation
The topic of this article is $q$-differential operators for $q$-spinor variables, the motivation comes from the study of the differential and integral calculus in $q$-spinor variables studied in [8] and some differential operator, in particular the Dirac operator mentioned in references [3], [4], [6]. In accordance with the above, our interest here is to relate the $q$-spinor variables with some differential operators, especially the $q$-Dirac operator, and its implications with differential calculus, integral and differential equations. The main objective of this work is, therefore, to study the $q$-differential operator on the $q$-spinor variables and its corresponding differential and integral calculus, and its differential equations. Solutions to the differential equation in $q$-spinor variables were also found.
This paper is organized as follows. We briefly recall the preliminaries will be used in this paper in Sect.2. The $q$-differential operators for $q$-spinor variables, the $q$-spinor chain rule, the new $q$-differential operator, the $q$-Dirac differential operator, and the integral formulas in $q$-spinor variables are then proposed in Sect. 3. Finally in the last Section the discussion and some suggestions for further work are presented.
Notation
Taking into account the Definition 1.5 the functions on the $q$ spinor variables we will denote by $\Psi(u^{\alpha}_{\dot{\beta}})$ for all $\alpha,\dot{\beta}=1,2$ , and thus we can write (38) and (39) of the form
$\dfrac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}=\dfrac{\Psi((qu)^{\alpha}_{\dot{\beta}})-q\Psi(u^{\alpha}_{\dot{\beta}})}{(qu)^{\alpha}_{\dot{\beta}}-qu^{\alpha}_{\dot{\beta}}},\quad(qu)^{\alpha}_{\dot{\beta}}\neq qu^{\alpha}_{\dot{\beta}}.$ (50)
2 $q$-Differential operators for $q$-spinor variables
The aim of this section is to define the $q$-differential operator for $q$- spinor variables. To begin, first we will mention an important rule of the $q$- spinor differential calculus, that is the q- spinor chain rule.
2.1 The $q$-spinor chain rule
Proposition 2.1.
Let us consider a $q$-spinor function on the form $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$. The $q$-spinor chain rule can be expressed as
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}.$ (51)
Proof.
Consider the following derivative
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))-q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))}{(qx_{\mu})-q(x_{\mu})},$ (52)
multiplying by $\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}$, we can rewrite (52) as
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))-q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}{(qx_{\mu})-q(x_{\mu})},$ (53)
in virtue of (38) and denoting $\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}{(qx_{\mu})-q(x_{\mu})}$ by $\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\mu}}$, finally we have
$\frac{\partial^{q}\Psi}{\partial^{q}x_{j}}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{j}},$ (54)
which completes the proof. ∎
Now, we will define a new $q$-differential operator for $q$-spinor variables over an orthonormal basis of $\mathds{R}^{n}$, different to usual Dirac operator and Cauchy - Riemann operator mentioned in [3], [4], [6], and [7].
2.2 The new $q$-differential operator for $q$-spinor variables
The motivation comes from the construction of the any diiferential operator that satisfy the property $D_{q}^{2}=-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}$ for all $\mu,\nu=1,2,...,n$ over an orthonormal basis of $\mathds{R}^{n}$.
Proposition 2.2.
Let $\left\{\mathrm{e}_{1},\mathrm{e}_{2},...,\mathrm{e}_{n}\right\}$ be an orthonormal basis of $\mathds{R}^{n}$. The $q$- Differential operator $D^{q}$
$D_{q}=\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}},$ (55)
is defined over $\mathds{R}^{n}$ under the follows relations
$\mathrm{e}^{2}_{\mu}\frac{\partial_{q}^{2}}{\partial_{q}x^{2}_{\nu}}+\mathrm{e}^{2}_{\nu}\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}$ $=\ -2\delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}},$ (56)
$\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}$ $=\ \delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}},\quad\mu,\nu,\alpha=1,2,...n.$ (57)
Proof.
The proof is based on the following observation: the square of (55) is equivalent to $-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}$. Therefore, we first compute $D^{2}_{q}$ resulting
$D^{2}_{q}$ $=\ \mathrm{e}^{2}_{\mu}\frac{\partial_{q}^{2}}{\partial_{q}x^{2}_{\nu}}+\mathrm{e}^{2}_{\nu}\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}},$ (58)
and substituting (56) and (57) into (58) we get
$D^{2}_{q}$ $=\ -2\delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}+\delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}},$ (59)
and for $\mu=\alpha$ finally we obtain $-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}$, which completes the proof. ∎
According to above, we can said that this operator acts over a $q$-spinor function on the form $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))$. Therefore (55) can be expressed as
$D_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))=\mathrm{e}_{\nu}\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\nu}}.$ (60)
and the derivatives $\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\mu}}$ and $\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\nu}}$ can be determined using the $q$-spinor chain rule defined by (51). We will consider the following example
Example 2.3.
Let $\Psi(x_{\mu},x_{\nu})=\exp{ix_{\mu}}$ ) and $u^{1}_{\dot{2}}=q^{2}x_{\mu}$. Now, to obtain the differential operator (55), first we write $\Psi(u^{1}_{\dot{2}})$ as $\exp{iu^{1}_{\dot{2}}}$ ). Later we apply (51)
$\frac{\partial_{q}\Psi}{\partial_{q}x_{\mu}}$ $=\ \frac{\partial_{q}}{\partial_{q}u^{1}_{\dot{2}}}(\exp{iu^{1}_{\dot{2}}})\frac{\partial_{q}}{\partial_{q}x_{\mu}}(q^{2}x_{\mu}),$
[…]
$=\ q^{2}\frac{\exp{iqu^{1}_{\dot{2}}}-q\exp{iqu^{1}_{\dot{2}}}}{[qu]^{1}_{\dot{2}}-q(u^{1}_{\dot{2}})},$
finally we apply (55) to obtain
$D_{q}\Psi(u^{1}_{\dot{2}})=\mathrm{e}_{\nu}q^{2}\frac{\exp{iqu^{1}_{\dot{2}}}-q\exp{iqu^{1}_{\dot{2}}}}{[qu]^{1}_{\dot{2}}-q(u^{1}_{\dot{2}})}.$
The expression (55) can be applied to functions that not depends of $u^{\alpha}_{\dot{\beta}}$, That is, it depends only on the variables $x_{\mu},x_{\nu}$. To obtain this result, it is necessary to define the $q$- derivatives in terms of the $q$-spinor variable $x^{\alpha}$ and $x_{\dot{\beta}}$, and the elements $\mathrm{e}_{\mu},\mathrm{e}_{\nu}$ in the following definition
Definition 2.4.
Let us consider the function $\psi(x_{\mu},x_{\nu})$. The expressions:
$\frac{\partial_{q}\psi}{\partial_{q}x_{\mu}}$ $=\ \frac{\psi(x_{\mu}+q\mathrm{e}_{\mu}x^{\alpha})-\psi(x_{\mu})}{x^{\alpha}},$ (61)
$\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}$ $=\ \frac{\psi(x_{\nu}+q\mathrm{e}_{\nu}x_{\dot{\beta}})-\psi(x_{\nu})}{x_{\dot{\beta}}},$ (62)
are the $q$-derivatives respect to $x_{\mu}$ and $x_{\nu}$ in terms of $q$- spinor variables $x_{\dot{\beta}},x^{\alpha}$ and the elements $\mathrm{e}_{\mu},\mathrm{e}_{\nu}$.
Let us consider the following remark.
Remark 2.5.
Taking into account the above claim, we can said that the differential operator not apply only over the functions $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$, but also over the function $\psi(x_{\mu},x_{\nu})$, and therefore the operator (55) also depends of the variables $x_{\mu}$ and $x_{\nu}$.
Example 2.6.
Consider the following function $\psi(x_{\mu},x_{\nu})=qx_{\nu}x_{\dot{\beta}},\dot{\beta}=\dot{2}$. Now, to determine the derivatives, we apply (61) and (62)
$\frac{\partial_{q}\psi}{\partial_{q}x_{\mu}}$ $=0$
$\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}$ $=\ \frac{q(x_{\nu}+q\mathrm{e}_{\nu}x_{\dot{2}})x_{\dot{2}}-qx_{\nu}x_{\dot{2}}}{x_{\dot{2}}}$
$=\ q^{2}\mathrm{e}_{\nu}x_{\dot{2}},$
thus the operator (55) is expressed as
$D_{q}\psi=q^{2}x_{\dot{2}}\mathrm{e}_{\mu}\mathrm{e}_{\nu}.$
2.3 The $q$-Dirac differential operator
Definition 2.7.
The $q$-analogue of (42) can be expressed as
$D_{\mu}^{q}=\gamma_{\mu}\frac{\partial^{q}}{\partial^{q}x_{\mu}}.$ (63)
We can now formulate our main results in the following propositions.
2.4 The $q$-differential operators for $q$-spinor variables
Proposition 2.8.
Let $\Psi$ be a function on the $q$-spinor variables. The operator (55) for $q$-spinor variables $D^{q}\Psi$ is given by
$D^{q}\Psi=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}u^{\alpha}_{\dot{\beta}}.$ (64)
Proof.
Let us consider the expressions (55) and
$\frac{\partial^{q}\Psi}{\partial^{q}x}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}}.$ (65)
Multiplying the left-hand side by $\mathrm{e}_{\nu}$ in (54), we have
$\begin{split}\mathrm{e}_{\nu}\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}&=\ \mathrm{e}_{\nu}\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}\\ &=\ \frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\mathrm{e}_{\nu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}},\end{split}$ (66)
and multiplying the left-hand side by $\mathrm{e}_{\mu}$ in (65) we get
$\begin{split}\mathrm{e}_{\mu}\frac{\partial^{q}\Psi}{\partial^{q}x_{\nu}}&=\ \mathrm{e}_{\mu}\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}}\\ &=\ \frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\mathrm{e}_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}},\end{split}$ (67)
adding (66) with (67) and considering (55), finally results
$D^{q}\Psi=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}u^{\alpha}_{\dot{\beta}},$
which our assertion. ∎
Remark 2.9.
The above proof implies the following relation
$\mathrm{e}_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}}-\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}}\mathrm{e}_{\mu}=0.$ (68)
Remark 2.10.
If $\Psi$ also depends of $x_{\nu}$, then
$D^{q}\Psi=\frac{\partial_{q}\Psi}{\partial^{q}x_{\nu}}D^{q}x_{\nu}.$ (69)
Remark 2.11.
The $q$-differential for the coordinate $x_{\mu}$ is given by
$D^{q}x_{\nu}:=\mathrm{e}_{\mu}\mathrm{d}^{q}x_{\nu},$ (70)
consequently the $q$-differential for $u^{\alpha}_{\dot{\beta}}$ is defined as
$D^{q}u^{\alpha}_{\dot{\beta}}:=\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\nu}}D^{q}x_{\nu}.$ (71)
Proposition 2.12.
Let $\Psi$ be a function on the $q$-spinor variables. The Dirac operator for $q$-spinor variables $D_{\mu}^{q}\Psi$ is given by
$D^{q}_{\mu}\Psi=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}.$ (72)
Proof.
Multiplying the left-hand side by $\gamma_{\mu}$ in (54), we have
$\begin{split}\gamma_{\mu}\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}&=\ \gamma_{\mu}\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}\\ &=\ \frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\gamma_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}},\end{split}$ (73)
and considering (63), finally we get
$D_{\mu}^{q}\Psi=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}_{\mu}u^{\alpha}_{\dot{\beta}},$
which is our claim. ∎
Remark 2.13.
The above proof implies the following relation
$\gamma_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}-\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}\gamma_{\mu}=0.$ (74)
Remark 2.14.
The Dirac $q$-differential for the coordinate $x_{\mu}$ is given by
$D^{q}_{\mu}x:=\gamma_{\mu}\mathrm{d}^{q}x^{\mu},$ (75)
consequently the Dirac $q$-differential for $u^{\alpha}_{\dot{\beta}}$ is defined as
$D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}:=\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x^{\mu}}D^{q}_{\mu}x^{\mu}.$ (76)
Now, by (52), (64) and (72), we have the follows $q$-spinor integral formulas showed in the following subsection.
2.5 The integral formulas in $q$-spinor variables
Proposition 2.15.
Let $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$ be a $q$-spinor function, and let $\Gamma_{q}$ be the closed contour of the deformed quantum complex plane, and $x_{0}\in\Gamma_{q}$. The integral formulas of the $q$-spinor variables are given by
$\oint_{\Gamma_{q}}\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})}$ $=\ \sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(qu^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n},$ (77)
$\oint_{\Gamma_{q}}\frac{\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{qu^{\alpha}_{\dot{\beta}}(x_{\mu})-(qu)^{\alpha}_{\dot{\beta}}(x_{0})}$ $=\ \frac{1}{q}\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n},$ (78)
where $\gamma_{\mu}$ are the $q$-deformed Dirac matrices defined in (43).
Proof.
The proof of this result is based on the ideas presented in [8]. Combining (38) with (72) we get
$D^{q}_{\mu}\Psi=\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}-qu^{\alpha}_{\dot{\beta}}}-\frac{q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}-qu^{\alpha}_{\dot{\beta}}},$ (79)
now, we integrate over the closed contour $\Gamma_{q}$ and we take $x_{0}\in\Gamma_{q}$ to obtain
$\oint_{\Gamma_{q}}D^{q}_{\mu}\Psi=\oint_{\Gamma_{q}}\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})}-\oint_{\Gamma_{q}}\frac{q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})}.$ (80)
Hence, to solve the integral $\oint_{\Gamma_{q}}D^{q}_{\mu}\Psi$, we will use similary the proof of the Theorem 2.9 of the reference [8], interchanging the Pauli matrices of $q$-deformed Minkowski space (e.g. [14]) by the $q$-deformed Dirac matrices (43), obtaining
$\begin{split}\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}-\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}&=\\ \oint_{\Gamma_{q}}\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})}-\oint_{\Gamma_{q}}\frac{q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})},\end{split}$ (81)
and finally, equalating terms that depend on $\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu})$ and $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$ we obtain (77) and (78), and the proof is complete. ∎
We will mention an important consequence of the Definition 3.4. starting of Eqs. (77) and (78): the formulation of the integral $\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}}(x))d^{q}x_{\mu}$, which we will mention in the following theorem.
Theorem 2.16.
Let $\Gamma_{q}$ be a closed contour and suppose that $x_{0}\in\Gamma_{q}$. Then for a function on $q-$ spinor variables $\Psi(u^{\alpha}_{\dot{\beta}}(x))$ the $q$-spinor integral formula is given by
$\frac{1}{qb}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu},\quad b\neq 0.$ (82)
Proof.
To formulate we can consider the following differential equation in $q$- spinor variables
$D^{q}_{\mu}\Psi(u^{\alpha}_{\dot{\beta}})-b\Psi(u^{\alpha}_{\dot{\beta}})=0.$ (83)
Multiplying on both sides of (83) by $\mathrm{d}^{q}x_{\mu}$, and using (63) we get
$\begin{split}\frac{\partial^{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}\mathrm{d}^{q}x_{\mu}&=\ b\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu}\\ \frac{\partial^{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\gamma_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\mu}}\mathrm{d}^{q}x_{\mu}&=\ b\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu},\end{split}$ (84)
taking into account (74), (75) and (76), we can rewrite (84) as
$\frac{\partial^{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D_{\mu}^{q}u^{\alpha}_{\dot{\beta}}=b\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu},$ (85)
from (38) it follows that
$\frac{\partial^{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial^{q}u^{\alpha}_{\dot{\beta}}}=\dfrac{q\Psi(u^{\alpha}_{\dot{\beta}})}{qu^{\alpha}_{\dot{\beta}}-(qu)^{\alpha}_{\dot{\beta}}},$ (86)
substituting (86) into (85) gives
$\left[\frac{q\Psi(u^{\alpha}_{\dot{\beta}})}{qu^{\alpha}_{\dot{\beta}}-(qu)^{\alpha}_{\dot{\beta}}}\right]D_{\mu}^{q}u^{\alpha}_{\dot{\beta}}=b\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu},$ (87)
we continue in this fashion integrating over $\Gamma_{q}$ on both sides for $x_{0}\in\Gamma_{q}$ , and using (78), finally we get
$\frac{1}{qb}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu},\quad b\neq 0,$
which is our claim. This expression is called the $q$-spinor integral formula for $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$. ∎
The same reasoning applies to the differential equation $D^{q}_{\mu}\Psi((qu)^{\alpha}_{\dot{\beta}})-b\Psi((qu)^{\alpha}_{\dot{\beta}})=0$ to obtain the integral in $q$-spinor variables
$\frac{1}{qb}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\oint_{\Gamma_{q}}\Psi((qu)^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x_{\mu}.$ (88)
Remark 2.17.
The expression (82) also can be expressed in virtue of (75) as
$\frac{\gamma^{\mu}}{qb}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}})D^{\mu}_{q}x.$ (89)
Notice that the expression (82) is not implies the final solution of (83).
Remark 2.18.
Similar arguments apply to the new $q$-differential operator (55), resulting
$\frac{\mathrm{e}^{\mu}}{qb}\left\{\sum\limits_{n=0}^{\infty}\left[\mathrm{e}_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}})D^{q}_{\mu}x.$ (90)
Proof.
It is sufficient to replace $\gamma_{\mu}$ by $\mathrm{e}_{\mu}$, to obtain (90), considering the differential equation in $q$-spinor variables of the form
$D^{q}\Psi(u^{\alpha}_{\dot{\beta}})-b\Psi(u^{\alpha}_{\dot{\beta}})=0,$ (91)
where $D^{q}$ is the new $q$-differential operator given by (55). ∎
However we will can solve differential equations in $q$-spinor variables of the form
$D^{q}_{\mu}\psi(u^{\alpha}_{\dot{\beta}})-b\phi(u^{\alpha}_{\dot{\beta}})$ = 0,absent 0=\ 0,= 0 , (92)
$D^{q}\psi(u^{\alpha}_{\dot{\beta}})-a\phi(u^{\alpha}_{\dot{\beta}})$ = 0absent 0=\ 0= 0 (93)
which we will show in the following section.
3 Differential equations in $q$-spinor variables
In order to obtain the solution of (92), it is necessary to put the following condition on $\psi$
$\oint_{\Gamma_{q}}D_{\mu}^{q}\psi(u^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x^{\mu}=\psi(u^{\alpha}_{\dot{\beta}}(x_{0})),\quad x_{0}\in\Gamma_{q}.$ (94)
Therefore, integrating both sides with respect to $x^{\mu}$ in (92) , applying (94), we get
$\begin{split}\oint_{\Gamma_{q}}D^{q}_{\mu}\psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu}&=\ b\oint_{\Gamma_{q}}\phi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu},\\ \psi(u^{\alpha}_{\dot{\beta}}(x_{0}))&=\ b\oint_{\Gamma_{q}}\phi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu},\\ \psi(u^{\alpha}_{\dot{\beta}}(x_{0}))&=\ b\oint_{\Gamma_{q}}\phi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu}\end{split}$ (95)
and applying (82) we obtain
$\psi(u^{\alpha}_{\dot{\beta}}(x_{0}))=\frac{1}{q}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\phi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}.$ (96)
Lemma 3.1.
We can generalize the solution (96) for all $x\in\Gamma_{q}$ as
$\psi(u^{\alpha}_{\dot{\beta}}(x))=\frac{1}{q}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\phi(u^{\alpha}_{\dot{\beta}}(x))\right]^{n}\right\}.$ (97)
To solve (93) we will consider the following remarks
Remark 3.2.
We first consider the new $q$-differential operator given by (55). Now, to solve the integral $\oint_{\Gamma_{q}}D^{q}\Psi$, we will use similary the proof of the Theorem 2.9 of the reference [8] (also can see the proof of Proposition 3.10.) to obtain similar expressions to (96) and (97), i.e.
$\oint_{\Gamma_{q}}D^{q}\psi(u^{\alpha}_{\dot{\beta}}(x_{0}))$ $=\ \psi(u^{\alpha}_{\dot{\beta}}(x_{0})),$ (98)
$\psi(u^{\alpha}_{\dot{\beta}}(x_{0}))$ $=\ \frac{1}{q}\left\{\sum\limits_{n=0}^{\infty}\left[\mathrm{e}_{\mu}\phi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\},$ (99)
Now, we will consider the following examples
Example 3.3.
Consider the differential equation in $q$-spinor variables of the form
$D^{q}_{\mu}\Psi(u^{\alpha}_{\dot{\beta}})-e\gamma^{\mu}A^{q}_{\mu}(x)\Psi(u^{\alpha}_{\dot{\beta}})-mg(u^{\alpha}_{\dot{\beta}})=0,\quad e\in\mathds{R},$ (100)
being $A^{q}_{\mu}(x)$ a $q$-potential function. Now, to solve (100), we can proceed analogously to the solution of (92) applying (94)
$\begin{split}\oint_{\Gamma_{q}}D^{q}_{\mu}\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu}-e\gamma^{\mu}\oint_{\Gamma_{q}}A^{q}_{\mu}(x)\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu}&=\ m\oint_{\Gamma_{q}}g(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu},\\ \Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))-e\gamma^{\mu}\oint_{\Gamma_{q}}A^{q}_{\mu}(x)\Psi(u^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x^{\mu}&=\ m\oint_{\Gamma_{q}}g(u^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x^{\mu},\end{split}$ (101)
and using (82) we obtain finally
$\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))+\frac{e}{qm}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}A^{q}_{\mu}(x_{0})\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\frac{1}{q}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}g(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}.$ (102)
Now, let us see other example.
Example 3.4.
Consider the differential equation in $q$-spinor variables (similar to (100)) of the form
$\gamma^{\mu}\partial^{q}_{\mu}\Psi(u^{\alpha}_{\dot{\beta}})-e\gamma^{\mu}A^{q}_{\mu}(x)\Psi(u^{\alpha}_{\dot{\beta}})-m\Psi(u^{\alpha}_{\dot{\beta}})=0,\quad e\in\mathds{R},$ (103)
where $A^{q}_{\mu}(x)$ is the same $q$-potential function of above example. This follows by the same method as in the above example, obtaining
$\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))+\frac{e}{qm}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}A^{q}_{\mu}(x_{0})\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\frac{1}{q}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}.$ (104)
Remark 3.5.
The expression (103) can be written as
$(\gamma^{\mu}\mathcal{D}^{q}_{\mu}-m)\Psi(u^{\alpha}_{\dot{\beta}})=0,$ (105)
where $\mathcal{D}^{q}_{\mu}=\partial^{q}_{\mu}-eA^{q}_{\mu}(x)$ is the q- covariant derivative.
Example 3.6.
Let us consider the differential equation in $q$- spinor variables $aD^{q}\psi+b\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x))=0$, where $B^{\mu}_{q}(x)$ is some $q$-arbitrary potential function. Therefore
$D^{q}\psi=-\frac{b}{a}\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x)),$ (106)
Using (55) (only the second contribution) we get
$\mathrm{e}_{\mu}\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}$ $=\ -\frac{b}{a}\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x)),$
multiplying both sides by $\mathrm{d}^{q}x_{\nu}$ results and applying (70)
$\mathrm{e}_{\mu}\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}\mathrm{d}^{q}x_{\nu}$ $=\ -\frac{b}{a}\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x_{\nu}$
$\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}D^{q}x_{\nu}$ $=\ -\frac{b}{a}\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x_{\nu}$
$D^{q}\psi$ $=\ -\frac{b}{a}\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x_{\nu},$
integrating over the closed contour $\Gamma_{q}$ , considering $x_{0}\subset\Gamma_{q}$ and using (90), (98), (99) and (69) into (106) we get
$\psi(u^{\alpha}_{\dot{\beta}}(x_{0}))=-\frac{1}{aq}\mathrm{e}^{\mu}\left\{\sum\limits_{n=0}^{\infty}\left[\mathrm{e}_{\mu}B^{\mu}_{q}(x_{0})\phi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}.$ (107)
and for all $x\in\Gamma_{q}$
$\psi(u^{\alpha}_{\dot{\beta}}(x))=-\frac{1}{aq}\mathrm{e}^{\mu}\left\{\sum\limits_{n=0}^{\infty}\left[\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x))\right]^{n}\right\}.$
4 Discussion and suggestions for further work
In Section 2, the equations (51), (60) and (72) describe some $q$-differential operators for $q$-spinor variables. Respect to (51), we can said that the function on the $q$-spinor variables does not depend on ly on the variable $u^{\alpha}_{\dot{\beta}}$ but also on $x_{\nu}$. From this result, the new $q$- differential operator for $q$-spinor variables expressed by (60) was proposed. This operator is motivated from the construction of the any differential operator that satisfy the property $D_{q}^{2}=-\frac{\partial^{2}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x_{\nu}^{2}}$ for all $\mu,\nu=1,2,...,n$. To obtain $D^{2}_{q}$ it is necessary to use the relations (56) and (57). This operator is different to usual Dirac operator and Cauchy Riemann operator mentioned in [3], [4], [6], and [7]. In the case of the Dirac operator for $q$-spinor variables, defined by (72), has been stablished from the Remarks 2.10 and 2.11. This operator is expressed in terms of the $q$-deformed Dirac matrices mentioned by Schmidt [15]. From the $q$-deformed Dirac operator, we define the integral formulas in $q$-spinor variables with the aim to solve the differential equations in $q$-spinor variables. Physically we can said that the Example 3.4 describes the Dirac equation for the electromagnetic case on the $q$-spinor variables, and furthermore the potential $A^{q}_{\mu}(x)$ can be interpreted as the q- electromagnetic potential. There are two further topics arising from this paper which are worth investigation., there is the problem of describing the Maxwell Electrodynamic Algebra which is defined by the following commutation relations
$A_{\mu}A^{\mu}$ $=\ |A_{0}|^{2}-A_{X}^{2},\quad X=1,2,3,$ (108)
$f_{\mu\nu}$ $=\ \partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},$ (109)
$\partial_{\mu}A^{\mu}=\partial_{\nu}A^{\nu}$ = 0,absent 0=\ 0,= 0 , (110)
$D_{\mu}$ $=\ \partial_{\mu}-eA_{\mu},\quad e\in\mathds{R}$ (111)
$\partial^{\mu}f_{\mu\nu}$ $=\ \Gamma_{\nu},$ (112)
$\partial_{\nu}f_{0}^{\mu\nu}$ = 0,absent 0=\ 0,= 0 , (113)
where $f_{0}^{\mu\nu}=\varepsilon^{\mu\nu 0}f_{\mu\nu}$ , $f^{\mu\nu}=0$ if $\mu=\nu$ and $f^{\mu\nu}\neq 0$ in otherwise. Finally, from (103) and taking into account the above, one can propose the $q-$ Dirac - Maxwell algebra, which is subject to relations
$f^{q}_{\mu\nu}$ $=\ \mathcal{D}^{q}_{\mu}A^{q}_{\nu}-\mathcal{D}^{q}_{\nu}A^{q}_{\mu},$ (114)
$\textit{{D}}^{q}$ $=\ \boldsymbol{\partial}^{q}-e\textit{{A}}\quad e\in\mathds{R},$ (115)
being $\textit{{D}}^{q}=\boldsymbol{\gamma}^{\mu}\mathcal{D}^{q}_{\mu},\boldsymbol{\partial}^{q}=\boldsymbol{\gamma}^{\mu}\partial^{q}_{\mu}$ and $\textbf{e}=\gamma^{\mu}A_{\mu}$, where $\boldsymbol{\gamma}^{\mu},\mu=1\cdots n$ are the generators for the Clifford algebras, and (115) is called the Covariant Derivative. Other suggestion is the formulation of the $q-$ Real Spinor Calculus based on the work of Zatloukal [16], which is defined by the following expressions for the derivatives
$\frac{\boldsymbol{\partial}_{q}\psi}{\boldsymbol{\partial}_{q}\textit{{x}}_{\dot{\alpha}}^{\beta}}=\frac{\psi(\textit{{x}}^{\beta}_{\dot{\alpha}}+q\textit{{u}}^{\beta}_{\dot{\alpha}})-q\psi(\textit{{u}}^{\beta}_{\dot{\alpha}})}{\textit{{x}}^{\beta}_{\dot{\alpha}}},$ (116)
where $\textit{{u}}^{\beta}_{\dot{\alpha}}=(\gamma_{\mu}\gamma_{\nu}u)^{\beta}_{\dot{\alpha}}$. The chain rule
$\frac{\boldsymbol{\partial}_{q}\Psi}{\boldsymbol{\partial}_{q}x_{j}}=\frac{\boldsymbol{\partial}_{q}\Psi}{\boldsymbol{\partial}_{q}\textit{{x}}^{\beta}_{\dot{\alpha}}}\frac{\boldsymbol{\partial}_{q}\textit{{x}}^{\beta}_{\dot{\alpha}}}{\boldsymbol{\partial}_{q}x_{j}},$ (117)
the $q$-difference operator for $q-$ real spinor variables
$\textit{{D}}_{2}^{q}$ $=\ \hat{\gamma}_{2}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}x_{2}},$ (118)
$\textit{{D}}^{q}_{j}$ $=\ i\hat{\gamma}_{5}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}x_{j}},$ (119)
$\underline{\textit{{D}}}_{j}^{q}$ $=\ i\hat{\gamma}_{2}\hat{\gamma}_{5}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}x_{j}},\quad j=1,..,5.$ (120)
For a function $\psi:(\textit{{v}}_{\dot{k}},\textit{{p}}_{\dot{k}})\longrightarrow\mathds{R}^{m}$, the $q$-conjugated derivatives are defined as
$\frac{\boldsymbol{\partial}_{q}\psi}{\boldsymbol{\partial}_{q}\textit{{v}}_{\dot{k}}}$ $=\ \frac{\psi(\textit{{v}}_{\dot{k}}+q\textit{{x}}^{\beta}_{\dot{\alpha}})-q\psi(\textit{{x}}^{\beta}_{\dot{\alpha}})}{\textit{{v}}_{\dot{k}}},$ (121)
$\frac{\boldsymbol{\partial}_{q}\psi}{\boldsymbol{\partial}_{q}\textit{{p}}_{\dot{k}}}$ $=\ \frac{\psi(\textit{{p}}_{\dot{k}}+q\textit{{u}}^{\beta}_{\dot{\alpha}})-q\psi(\textit{{u}}^{\beta}_{\dot{\alpha}})}{\textit{{p}}_{\dot{k}}}.$ (122)
The $q$-difference operators associated to conjugated real spinor variables are given by
$\textit{{D}}_{q}$ $=\ \frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}\textit{{v}}_{\dot{0}}}+\gamma_{1}\gamma_{3}\frac{\boldsymbol{\partial_{q}}}{\boldsymbol{\partial}_{q}\boldsymbol{v}_{\dot{1}}}+i\gamma_{3}\gamma_{\dot{0}}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}\textit{{v}}_{\dot{2}}}+\gamma_{1}\gamma_{2}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}\textit{{v}}_{\dot{3}}},$ (123)
$\textit{{D}}^{\prime}_{q}$ $=\ \frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}\textit{{p}}_{\dot{0}}}+\gamma_{1}\gamma_{3}\frac{\boldsymbol{\partial_{q}}}{\boldsymbol{\partial}_{q}\textit{{p}}_{\dot{1}}}+i\gamma_{3}\gamma_{0}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}\textit{{p}}_{\dot{2}}}+\gamma_{1}\gamma_{2}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}\textit{{p}}_{\dot{3}}},$ (124)
and the $q$-spinor real integral formulas of the $q$-spinor conjugated variables are given by
$\int_{\Omega_{q}}\frac{\psi(q\textit{{v}}_{\dot{k}})\mathbf{d}_{q}\textit{{v}}_{\dot{k}}}{\textit{{v}}_{\dot{k}}-\textit{{x}}^{\beta}_{\dot{\alpha}}}$ $=\ q\sum\limits_{n=0}^{\infty}[\gamma^{\mu}\gamma^{\nu}\psi(q\textit{{x}}^{\beta}_{\dot{\alpha}})]^{n},$ (125)
$\int_{\Omega_{q}}\frac{\psi[(1-q^{-1})\textit{{v}}_{\dot{k}}]\mathbf{d}_{q}\textit{{v}}_{\dot{k}}}{\textit{{v}}_{\dot{k}}-\textit{{x}}^{\beta}_{\dot{\alpha}}}$ $=\ \sum\limits_{n=0}^{\infty}[\gamma^{\mu}\gamma^{\nu}\psi[(1-q^{-1})\textit{{x}}^{\beta}_{\dot{\alpha}}]]^{n},$ (126)
$\int_{\Omega_{q}}\frac{\psi(q\textit{{p}}_{\dot{k}})\mathbf{d}_{q}\textit{{p}}_{\dot{k}}}{\textit{{p}}_{\dot{k}}-\textit{{u}}^{\beta}_{\dot{\alpha}}}$ $=\ q\sum\limits_{n=0}^{\infty}[\gamma^{\mu}\gamma^{\nu}\psi(q\textit{{u}}^{\beta}_{\dot{\alpha}})]^{n},$ (127)
$\int_{\Omega_{q}}\frac{\psi[(1-q^{-1})\textit{{p}}_{\dot{k}}]\mathbf{d}_{q}\textit{{p}}_{\dot{k}}}{\textit{{p}}_{\dot{k}}-\textit{{u}}^{\beta}_{\dot{\alpha}}}$ $=\ \sum\limits_{n=0}^{\infty}[\gamma^{\mu}\gamma^{\nu}\psi[(1-q^{-1})\textit{{u}}^{\beta}_{\dot{\alpha}}]]^{n}.$ (128)

### v2

@@
 Abstract
-In this paper we introduce the $q$-differential operator for $q$-spinor variables.
-We establish the $q$-spinor chain rule , the new $q$-differential operator, the $q$-Dirac differential operators and the $q$-complex spinor integrals.
-We also define the $q$-spinor differential equation.
-The suggestions for further work at the end of the paper.
+We introduce a q-differential operator adapted to q-spinor variables, establishing a corresponding q-spinor chain rule and defining both standard and Dirac-type q-differential operators.
+Integral formulas in q-spinor variables are derived, and applications to q-deformed spinor differential equations are explored through explicit examples.
+The framework extends existing q-calculus to spinorial structures, offering potential insights into quantum deformations of relativistic field equations.
+We conclude with suggestions for future developments, including a q-analogue of the Dirac–Maxwell algebra.
 Keywords:
@@
 1 Introduction
-Based on the work of Beretetskii et al., Lachieze-Rey, Gori et al., and Cartan, the spinor $\psi^{\alpha}$ is defined as a magnitude components $\alpha=1,2$ expressed as $\psi^{\alpha}=\begin{bmatrix}\psi_{1}\\ \psi_{2}\end{bmatrix}$ and its complex conjugate $\dot{\varphi}$ in terms of the rotation matrices (see [2],[1],[9],[5] for more details).
-Based on the work of the previously mentioned authors, there are two types of operations on spinors, which are reflections and rotations.
-In group theory, the set of rotations described by the matrices with complex entries is group $SU(2)$, whose generators are the Pauli matrices, described in the work of Zettili [17].
-With respect to rotation matrices, Gori et al.
-mention, in their work, the rotation matrices that originated the Pauli matrices in the form:
-$R_{x}(\theta)=\begin{bmatrix}\cos(\theta/2)&i\sin(\theta/2)\\ i\sin(\theta/2)&\cos(\theta/2)\end{bmatrix},R_{y}(\theta)=\begin{bmatrix}\cos(\theta/2)&\sin(\theta/2)\\ \sin(\theta/2)&\cos(\theta/2)\end{bmatrix},R_{z}(\theta)=\begin{bmatrix}e^{i\theta/2}&0\\ 0&e^{-i\theta/2}\end{bmatrix},$ (1)
-being $\theta$ the angle of rotation [5].
-Beretetskii et al.
-define the covariance and contravariance over the spinors by the relation $\psi^{\prime 1}=\psi_{2},\psi^{\prime 2}=-\psi_{1}$ from the matrix $g_{\alpha\beta}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ and, similarly, for pointed spinors, $\psi_{\dot{1}}=\psi^{\dot{2}},\psi_{\dot{2}}=-\psi^{\dot{1}}$ [1].
-The same author defines bispinors as the pair $(\psi^{\alpha},\varphi_{\dot{\alpha}})$, which form a broader group of Lorentz, and, with them, the scalar product is formulated as $(\psi^{\alpha},\varphi_{\dot{\alpha}})\cdot(f^{\alpha},h_{\dot{\alpha}})$.
-The author of the reference [9] mentions algebra $Cl(3)$ as a space-time formulation generated by vectors $e_{\mu}$, which form a basis for $\mathds{R}^{1,3}$ that satisfies the relation $e_{\mu}\cdot e_{\nu}=g_{\mu\nu}$, inducing a 16 -membered basis, as described below:
-  • 1
-    $1$ scalar,
-  • 2
-    $(e_{0},e_{1},e_{2},e_{3})$ 4-vector,
-  • 3
-    $(e_{0}e_{1},e_{0}e_{2},e_{0}e_{3},e_{1}e_{2},e_{2}e_{3},e_{3}e_{1})$ 6- bivectors,
-  • 4
-    $(e_{1}e_{2}e_{3},e_{0}e_{2}e_{3},e_{0}e_{1}e_{3},e_{0}e_{1}e_{2})$ 4-trivectors,
-  • 5
-    $e_{5}\equiv e_{0}e_{1}e_{2}e_{3}$ pseudoescalar.
-In accordance with the above, the same author describes the Weyl spinors as
-$\begin{bmatrix}\psi^{1}\\ \psi^{2}\end{bmatrix}$ and the Dirac spinor $\begin{bmatrix}\psi^{\alpha}\\ \varphi^{\dot{\alpha}}\end{bmatrix}$, as defined in the work of Beretetskii et al [1].
-The $q$ Lorentzian algebra was defined in the reference [12].
-The quantum complex spinors have components $\psi_{1}$ and $\psi^{2}$ and conjugates $\varphi^{\dot{1}}$ and $\varphi_{\dot{2}}$.
-For all $q\in\mathds{R}-\left\{0\right\}$, they satisfy the following $q$ relations
-$\psi_{1}\psi^{2}=q\psi^{2}\psi_{1},$ $\ \psi^{2}\varphi^{\dot{1}}=q\varphi^{\dot{1}}\psi^{2},$ (2)
-$\psi_{1}\varphi^{\dot{1}}=\varphi^{\dot{1}}\psi_{1}-q(q+q^{-1})^{1/2}\varphi_{\dot{2}}\varphi^{\dot{1}},$ $\ \psi^{2}\varphi^{\dot{1}}=\varphi^{\dot{1}}\psi^{2},$ (3)
-$\psi_{1}\varphi_{\dot{2}}=q\varphi_{\dot{2}}\psi_{1},$ $\ \varphi^{\dot{1}}\varphi_{\dot{2}}=q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}}.$ (4)
-Considering spinors $\psi^{\alpha},\varphi^{\dot{\alpha}}$ and $\left\{\tau^{1},T^{2},S^{1},\sigma^{2}\right\}$ as the generators of the $q$ Lorentzian algebra for the group $U_{q}(su(2))$ [13], we have:
-  • 1.
-    For $\psi^{\alpha},\alpha=1,2$
-    $\tau^{1}\psi_{1}=$ $\ \psi_{1}\tau^{1},$ (5)
-    $\tau^{1}\psi^{2}=$ $\ \psi^{2}\tau^{1}-q(q+q^{-1})^{2}\psi_{1}T^{2},$ (6)
-    $T^{2}\psi_{1}=$ $\ q^{-1}\psi_{1}T^{2},$ (7)
-    $S^{1}\psi_{1}=$ $\ q\psi_{1}S^{1},$ (8)
-    $T^{2}\psi^{2}=$ $\ q\psi^{2}T^{2},$ (9)
-    $S^{1}\psi^{2}=$ $\ q^{-1}\psi^{2}S^{1}-\psi_{1}\sigma^{2},$ (10)
-    $\sigma^{2}\psi_{1}=$ $\ \psi_{1}\sigma^{2},$ (11)
-    $\sigma^{2}\psi^{2}=$ $\ \psi^{2}\sigma^{2},$ (12)
-  • 2.
-    Their complex conjugates $\varphi^{\dot{\alpha}},\alpha=\dot{1},\dot{2}$
-    $\tau^{1}\varphi^{\dot{1}}=$ $\ q^{-1}\varphi^{\dot{1}}\tau^{1},$ (13)
-    $\tau^{1}\varphi_{\dot{2}}=$ $\ q\varphi_{\dot{2}}\tau^{1},$ (14)
-    $T^{2}\varphi^{\dot{1}}=$ $\ \varphi^{\dot{1}}T^{2}+q^{-1}\varphi_{\dot{2}}\tau^{1},$ (15)
-    $T^{2}\varphi_{\dot{2}}=$ $\ \varphi_{\dot{2}}T^{2},$ (16)
-    $S^{1}\varphi^{\dot{1}}=$ $\ \varphi^{\dot{1}}S^{1},$ (17)
-    $\sigma^{2}\varphi^{\dot{1}}=$ $\ q\varphi^{\dot{1}}\sigma^{2}+(q+q^{-1})^{2}\varphi_{\dot{2}}S^{1},$ (18)
-    $\sigma^{2}\varphi_{\dot{2}}=$ $\ q\varphi_{\dot{2}}\sigma^{2}.$ (19)
-Deformed commutation relations for $q$- Lorentzian algebra are defined in the next proposition, on the quantum-symmetric plane and the quantum anti-symmetric plane.
-Proposition 1.1.
-Consider generator $T$ of the set $\left\{\tau^{1},T^{2},S^{1},\sigma^{2}\right\}$ for the algebra $U_{q}(su(2))$ and the relations 2, 3, and 4 , defined in [12] and [13] .
-The $q$ Lorentzian algebra for spinors in the deformed space is defined through the following relations:
-$T(\psi_{1}\psi^{2}-q\psi^{2}\psi_{1})=$ $\ (\psi_{1}\psi^{2}-q\psi^{2}\psi_{1})T,$ (20)
-$T(\psi_{1}\psi^{2}-q\psi^{2}\psi_{1})=$ $\ (\psi_{1}\psi^{2}-q\psi{2}\psi^{1})T,$ (21)
-$T(\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})=$ $\ (\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})T,$ (22)
-$T(\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})=$ $\ (\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})T.$ (23)
-Definition 1.2.
-The following are the bosonic $q$ deformed Minkowskian Pauli spin matrices defined in the Schmidt work [13]:
-$(\sigma^{+})_{\alpha\dot{\beta}}=\begin{bmatrix}0&0\\ 0&q\end{bmatrix},\quad(\sigma^{-})_{\alpha\dot{\beta}}=\begin{bmatrix}q&0\\ 0&0\end{bmatrix},$ (24)
-$(\sigma^{3})_{\alpha\dot{\beta}}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\sigma^{0})_{\alpha\dot{\beta}}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&-q^{-1/2}\\ q^{1/2}&0\end{bmatrix}.$ (25)
-Likewise, the conjugated Pauli matrices are:
-$(\overline{\sigma}^{+})_{\dot{\alpha}\beta}=\begin{bmatrix}0&0\\ 0&q^{-1}\end{bmatrix},\quad(\overline{\sigma}^{-})_{\dot{\alpha}\beta}=\begin{bmatrix}q^{-1}&0\\ 0&0\end{bmatrix},$ (26)
-$(\overline{\sigma}^{3})_{\dot{\alpha}\beta}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\overline{\sigma}^{0})_{\dot{\alpha}\beta}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{-1/2}\\ -q^{1/2}&0\end{bmatrix}.$ (27)
-The inverse Pauli matrices
-$(\sigma^{-1}_{+})_{\alpha\dot{\beta}}=\begin{bmatrix}0&0\\ 0&q^{-1}\end{bmatrix},\quad(\sigma^{-1}_{-})_{\alpha\dot{\beta}}=\begin{bmatrix}q^{-1}&0\\ 0&0\end{bmatrix},$ (28)
-$(\sigma^{-1}_{3})_{\alpha\dot{\beta}}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\sigma^{-1}_{0})_{\alpha\dot{\beta}}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&-q^{-1/2}\\ q^{1/2}&0\end{bmatrix}.$ (29)
-Finally
-$(\overline{\sigma}^{-1}_{+})_{\dot{\alpha}\beta}=\begin{bmatrix}0&0\\ 0&q^{-1}\end{bmatrix},\quad(\overline{\sigma}^{-1}_{-})_{\dot{\alpha}\beta}=\begin{bmatrix}q^{-1}&0\\ 0&0\end{bmatrix},$ (30)
-$(\overline{\sigma}^{-1}_{3})_{\dot{\alpha}\beta}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\overline{\sigma}^{-1}_{0})_{\dot{\alpha}\beta}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&-q^{-1/2}\\ q^{1/2}&0\end{bmatrix}.$ (31)
-Definition 1.3.
-The $q$ Lorentzian spinor variables or $q$ spinor variables are defined according to the expressions (2), (3), and (4) as follows:
-$u^{2}_{1}\equiv$ $\ \psi_{1}\psi^{2}-q\psi^{2}\psi_{1},$ (32)
-$v^{\dot{1}2}\equiv$ $\ \psi^{2}\varphi^{\dot{1}}-q\varphi^{\dot{1}}\psi^{2},$ (33)
-$x^{\dot{1}}_{1\dot{2}}\equiv$ $\ \psi_{1}\varphi^{\dot{1}}-\varphi^{\dot{1}}\psi_{1}+q(q+1)^{-1/2}\varphi_{\dot{2}}\varphi^{\dot{1}},$ (34)
-$y^{2\dot{1}}\equiv$ $\ \psi^{2}\varphi^{\dot{1}}-\varphi^{\dot{1}}\psi^{2},$ (35)
-$z^{\dot{1}}_{\dot{2}}\equiv$ $\ \psi_{1}\varphi_{\dot{2}}-q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}},$ (36)
-$t_{1\dot{2}}\equiv$ $\ \psi_{1}\varphi_{\dot{2}}-q\varphi_{\dot{2}}\psi_{1}.$ (37)
-Definition 1.4.
-We consider the set $U=\left\{u^{2}_{1},v^{\dot{1}2},x^{\dot{1}}_{1\dot{2}},z^{\dot{1}}_{\dot{2}},y^{2\dot{1}},t_{1\dot{2}}\right\}\subset\mathds{C}$.
-A function on the $q$ spinor variables is defined as $\Psi(U)=\Psi(u^{2}_{1},v^{\dot{1}2},x^{\dot{1}}_{1\dot{2}},z^{\dot{1}}_{\dot{2}},y^{2\dot{1}},t_{1\dot{2}})$.
-Definition 1.5.
-Let $f,g:U\longrightarrow\mathds{C}$ be functions and $u^{\beta}\in U$.
-The following properties are satisfy on the $q$ spinor variables, we state some clear properties of the functions on the $q$ spinor variables
-  • 1.
-    $(f+g)(u^{\beta})=f(u^{\beta})+g(u^{\beta})$.
-  • 2.
-    $(f\cdot g)(u^{\beta})=f(u^{\beta})\cdot g(u^{\beta})$.
-  • 3.
-    $(f-g)(u^{\beta})=f(u^{\beta})-g(u^{\beta})$.
-  • 4.
-    $\left(\dfrac{f}{g}\right)(u^{\beta})=\dfrac{f(u^{\beta})}{g(u^{\beta})},\quad g(u^{\beta})\neq 0$.
-Definition 1.6.
-For a function $f:U\longrightarrow\mathds{C}$ and $u^{\beta}\in\mathbb{C}$ , the $q$ spinor derivative is defined as [8] :
-$\dfrac{\mathrm{d}_{q}f}{\mathrm{d}_{q}u^{\beta}}=\dfrac{f((qu)^{\beta})-qf(u^{\beta})}{(qu)^{\beta}-qu^{\beta}},$ (38)
-and its conjugate complex
-$\dfrac{\mathrm{d}_{q}f}{\mathrm{d}_{q}v^{\dot{\alpha}}}=\dfrac{f((qv)^{\dot{\alpha}})-qf(v^{\dot{\alpha}})}{(qv)^{\dot{\alpha}}-qv^{\dot{\alpha}}}.$ (39)
-1.1 Clifford algebra and Dirac operator
-Let $\left\{\gamma_{1},\gamma_{2},\cdots,\gamma_{n},\right\}$ be an orthonormal basis of $\mathds{R}^{n}$. The Clifford algebra is generated over $\mathbb{R}^{n}$ under the relation
-$\gamma_{\mu}\gamma_{\nu}+\gamma_{\nu}\gamma_{\mu}=-2\delta_{\mu\nu}\gamma_{0},\quad\gamma_{\mu}^{2}=-|\gamma_{\mu}|^{2}\gamma_{0},\quad\mu,\nu=1,2,...,n,$ (40)
-where $\delta_{\mu\nu}$ is the Kronecker symbol (see [11], [3], [10] for more details).
-We will denote the Clifford algebra by $Cl_{n}$, and each element in $Cl_{n}$ can be expressed by its components as $\sum_{a}\gamma_{a}x_{a}$, where $a=(\mu_{1},...,\mu_{n})$ with each $\mu_{l}\in\left\{1,2,...,n\right\}$.
-Any element $\textit{{x}}\in\mathds{R}^{n}$ can be identified with a 1-vector in the Clifford algebra [10]
-$(x_{1},x_{2},...,x_{n})\longrightarrow\textit{{x}}=x_{1}\gamma_{1}+x_{2}\gamma_{2}+\cdots+x_{n}\gamma_{n}.$ (41)
-On other hand, the Dirac operator used here is
-$D:=\sum\limits_{\mu=1}^{n}\gamma_{\mu}\frac{\partial}{\partial x_{\mu}},$ (42)
-we refer to reader to [3], [6], [4] for more details.
-1.2 $q$-Deformed Dirac matrices
-Definition 1.7.
-The $q$-deformed Dirac matrices are defined in [15], and are given by
-$\gamma_{\mu}:=\ \begin{bmatrix}0&(\sigma_{\mu})^{\alpha}_{\dot{\beta}}\\ (\overline{\sigma}_{\mu})^{\dot{\alpha}}_{\beta}&0\end{bmatrix},$ (43)
-where $(\sigma_{\mu})^{\alpha}_{\dot{\beta}}$ and $(\overline{\sigma}_{\mu})^{\dot{\alpha}}_{\beta}$ denote the Pauli matrices of $q-$ deformed Minkowski space (e.g.
-[14] for more details), and are defined as
-$\begin{split}(\sigma_{+})^{\alpha}_{\dot{\beta}}=\begin{bmatrix}0&0\\ 0&kq^{1/2}\lambda^{1/2}_{+}\end{bmatrix},\quad(\sigma_{3})^{\alpha}_{\dot{\beta}}=k\begin{bmatrix}0&q\\ 1&0\end{bmatrix},\\ (\sigma_{-})^{\alpha}_{\dot{\beta}}=k\begin{bmatrix}q^{1/2}\lambda^{1/2}_{+}&0\\ 0&0\end{bmatrix},\quad(\sigma_{0})^{\alpha}_{\dot{\beta}}=k\begin{bmatrix}0&-q^{-1}\\ 1&0\end{bmatrix},\end{split}$ (44)
-and their conjugated counterparts
-$\begin{split}(\overline{\sigma}_{+})^{\alpha}_{\dot{\beta}}=\begin{bmatrix}0&0\\ 0&\overline{k}q^{-1/2}\lambda^{1/2}_{+}\end{bmatrix},\quad(\overline{\sigma}_{3})^{\alpha}_{\dot{\beta}}=\overline{k}\begin{bmatrix}0&1\\ q^{-1}&0\end{bmatrix},\\ (\overline{\sigma}_{-})^{\alpha}_{\dot{\beta}}=\overline{k}\begin{bmatrix}q^{-1/2}\lambda^{1/2}_{+}&0\\ 0&0\end{bmatrix},\quad(\overline{\sigma}_{0})^{\alpha}_{\dot{\beta}}=\overline{k}\begin{bmatrix}0&1\\ -q&0\end{bmatrix},\end{split}$ (45)
-where $k,\overline{k}$ are characteristic parameters associated to bosons ($q=+1$) and fermions ($q=-1$).
-1.3 $q$-Spinor complex integral formulas
-Definition 1.8.
-[8] Let $\Gamma_{q}$ be the closed contour of the deformed quantum complex plane, and $u_{0}^{\beta},v^{\dot{\alpha}}\subset\Gamma_{q}$ point spinors contained in the contour.
-The $q$- spinor complex integral formulas are defined by
-$\oint_{\Gamma_{q}}\dfrac{\Psi(u^{\beta})\mathrm{d}_{q}u^{\beta}}{(qu)^{\beta}-qu^{\beta}_{0}}$ $=\ \dfrac{1}{q}\sum_{n=0}^{\infty}\left[(\overline{\sigma}_{\mu})_{\dot{\alpha}\beta}\Psi(u_{0}^{\beta})\right]^{n},$ (46)
-$\oint_{\Gamma_{q}}\dfrac{\Psi((qu)^{\beta})\mathrm{d}_{q}u^{\beta}}{(qu)^{\beta}-qu^{\beta}_{0}}$ $=\ \sum_{n=0}^{\infty}\left[(\overline{\sigma}_{\mu})_{\dot{\alpha}\beta}\Psi((qu_{0})^{\beta})\right]^{n},$ (47)
-$\oint_{\Gamma_{q}}\dfrac{\Psi(v^{\dot{\alpha}})\mathrm{d}_{q}v^{\dot{\alpha}}}{(qv)^{\dot{\alpha}}-qv^{\dot{\alpha}}_{0}}$ $=\ \dfrac{1}{q}\sum_{m=0}^{\infty}\left[(\sigma_{\mu})_{\alpha\dot{\beta}}\Psi(v_{0}^{\dot{\alpha}})\right]^{m},$ (48)
-$\oint_{\Gamma_{q}}\dfrac{\Psi((qv)^{\dot{\alpha}})\mathrm{d}_{q}v^{\dot{\alpha}}}{(qv)^{\dot{\alpha}}-qv^{\dot{\alpha}}_{0}}$ $=\ \sum_{m=0}^{\infty}\left[(\sigma_{\mu})_{\alpha\dot{\beta}}\Psi((qv_{0})^{\dot{\alpha}})\right]^{m}.$ (49)
-Motivation
-The topic of this article is $q$-differential operators for $q$-spinor variables, the motivation comes from the study of the differential and integral calculus in $q$-spinor variables studied in [8] and some differential operator, in particular the Dirac operator mentioned in references [3], [4], [6].
-In accordance with the above, our interest here is to relate the $q$-spinor variables with some differential operators, especially the $q$-Dirac operator, and its implications with differential calculus, integral and differential equations.
-The main objective of this work is, therefore, to study the $q$-differential operator on the $q$-spinor variables and its corresponding differential and integral calculus, and its differential equations.
-Solutions to the differential equation in $q$-spinor variables were also found.
+The mathematical formulation of spinor fields is essential for the study of relativistic quantum mechanics, with foundations established in the works of Cartan [2], Berestetskii et al.
+[1], and others.
+Spinors are typically represented as two-component objects, such as $\psi^{\alpha}=\begin{bmatrix}\psi_{1}\\ \psi_{2}\end{bmatrix}$, and their conjugates are associated with transformation properties under the Lorentz group.
+The representation theory underlying these structures connects to the group $SU(2)$, whose generators are the Pauli matrices [17].
+Gori et al.
+[5] provide explicit matrix forms for rotations, revealing the origin of the Pauli matrices from representations of spinor transformations.
+Berestetskii et al.
+[1] formalised the covariant and contravariant behavior of spinors, as well as their scalar products and bispinor structures.
+Lachieze-Rey [10] further connected these ideas to Clifford algebras, particularly $Cl(3)$, which provides a geometric interpretation of space-time through multivector bases in $\mathds{R}^{1,3}$.
+In the context of quantum deformations, q-spinor variables and their associated algebraic relations emerge from the q-deformed Lorentz algebra [12].
+These variables satisfy non-trivial commutation relations, forming the foundation for defining q-differential operators.
+The structure of the q-Lorentzian algebra and its generators, such as $\tau^{1},T^{2},S^{1},\sigma^{2}$, governs the algebraic relations between spinors and their conjugates [13].
+Building on the formalism developed in [8], we define q-spinor variables and derive corresponding differential and integral operators, including q-derivatives and q-complex integrals.
+These operators act on functions defined over quantum spinor variables and respect the q-deformed algebraic relations.
+Moreover, the formulation introduces q-Pauli matrices and q-Dirac matrices [15], extending classical Clifford structures to a deformed setting.
+We also review the classical Dirac operator defined over Clifford algebras [11, 3, 4], and reinterpret its structure in terms of q-deformed analogues.
+The resulting q-Dirac operator inherits many key properties of its classical counterpart, including its role in defining and solving differential equations.
+The principal motivation of this work is to investigate the role of q-differential operators in spinor calculus, as introduced in [8], and to explore their potential in the formulation and solution of q-deformed differential equations.
+Inspired by earlier treatments of classical and quantum Dirac operators [6], our objective is to construct a consistent differential and integral calculus over q-spinor variables, develop explicit solutions to q-Dirac-type equations, and outline the implications of this theory for quantum-deformed field equations.
+We conclude the introduction by highlighting the long-term goal of formulating a q-Dirac–Maxwell algebra and a q-real spinor calculus, drawing from both geometric and algebraic perspectives [16].
+These efforts aim to enrich the interplay between quantum groups, spin geometry, and field theory, laying the groundwork for future developments in noncommutative quantum physics.
 This paper is organized as follows.
-We briefly recall the preliminaries will be used in this paper in Sect.2.
-The $q$-differential operators for $q$-spinor variables, the $q$-spinor chain rule, the new $q$-differential operator, the $q$-Dirac differential operator, and the integral formulas in $q$-spinor variables are then proposed in Sect.
-3.
+We briefly recall the preliminaries will be used in this paper in Section 2.
+The $q$-differential operators for $q$-spinor variables, the $q$-spinor chain rule, the new $q$-differential operator, the $q$-Dirac differential operator, and the integral formulas in $q$-spinor variables are then proposed in Section 3.
 Finally in the last Section the discussion and some suggestions for further work are presented.
-Notation
-Taking into account the Definition 1.5 the functions on the $q$ spinor variables we will denote by $\Psi(u^{\alpha}_{\dot{\beta}})$ for all $\alpha,\dot{\beta}=1,2$ , and thus we can write (38) and (39) of the form
-$\dfrac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}=\dfrac{\Psi((qu)^{\alpha}_{\dot{\beta}})-q\Psi(u^{\alpha}_{\dot{\beta}})}{(qu)^{\alpha}_{\dot{\beta}}-qu^{\alpha}_{\dot{\beta}}},\quad(qu)^{\alpha}_{\dot{\beta}}\neq qu^{\alpha}_{\dot{\beta}}.$ (50)
 2 $q$-Differential Operators for $q$-Spinor Variables
-The aim of this section is to define the $q$-differential operator for $q$- spinor variables.
-To begin, first we will mention an important rule of the $q$- spinor differential calculus, that is the q- spinor chain rule.
+The aim of this section is to define a $q$-differential operator for $q$-spinor variables.
+To begin with, we introduce a fundamental rule of the $q$-spinor differential calculus, namely the $q$-spinor chain rule.
 2.1 The $q$-Spinor Chain Rule
 Proposition 2.1.
-Let us consider a $q$-spinor function on the form $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$. The $q$-spinor chain rule can be expressed as
-$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}.$ (51)
+Let $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$ be a $q$-spinor function. Then the $q$-spinor chain rule is given by
+$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\cdot\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}.$ (1)
 Proof.
-Consider the following derivative
-$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))-q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))}{(qx_{\mu})-q(x_{\mu})},$ (52)
-multiplying by $\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}$, we can rewrite (52) as
-$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))-q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}{(qx_{\mu})-q(x_{\mu})},$ (53)
-in virtue of (38) and denoting $\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}{(qx_{\mu})-q(x_{\mu})}$ by $\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\mu}}$, finally we have
-$\frac{\partial^{q}\Psi}{\partial^{q}x_{j}}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{j}},$ (54)
+Consider the $q$-derivative of the composite function:
+$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\Psi\left((qu)^{\alpha}_{\dot{\beta}}(x_{\mu})\right)-q\,\Psi\left(u^{\alpha}_{\dot{\beta}}(x_{\mu})\right)}{qx_{\mu}-qx_{\mu}}.$ (2)
+Multiplying and dividing by the non-vanishing quantity $(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-q\,u^{\alpha}_{\dot{\beta}}(x_{\mu})$, we may rewrite (2) as
+$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\Psi\left((qu)^{\alpha}_{\dot{\beta}}(x_{\mu})\right)-q\,\Psi\left(u^{\alpha}_{\dot{\beta}}(x_{\mu})\right)}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-q\,u^{\alpha}_{\dot{\beta}}(x_{\mu})}\cdot\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-q\,u^{\alpha}_{\dot{\beta}}(x_{\mu})}{qx_{\mu}-qx_{\mu}}.$ (3)
+Denoting the first factor as $\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}$ and the second as $\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\mu}}$, we obtain
+$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\cdot\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\mu}},$ (4)
+which establishes the desired result. ∎
+We now proceed to define a novel $q$-differential operator acting on $q$-spinor variables over an orthonormal basis of $\mathds{R}^{n}$.
+This operator differs from the classical Dirac and Cauchy–Riemann operators considered in [3, 4, 6, 7, 9].
+2.2 The new $q$-differential operator for $q$-spinor variables
+The motivation arises from the construction of a differential operator satisfying
+$D_{q}^{2}=-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}},$
+for all $\mu,\nu=1,2,\dotsc,n$, over an orthonormal basis of $\mathds{R}^{n}$.
+Proposition 2.2.
+Let $\left\{\mathrm{e}_{1},\mathrm{e}_{2},\dotsc,\mathrm{e}_{n}\right\}$ be an orthonormal basis of $\mathds{R}^{n}$. The $q$-differential operator $D^{q}$ defined by
+$D_{q}=\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}},$ (5)
+satisfies the relations:
+$\mathrm{e}_{\mu}^{2}\frac{\partial_{q}^{2}}{\partial_{q}x_{\nu}^{2}}+\mathrm{e}_{\nu}^{2}\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}^{2}}$ $=-2\delta_{\mu\alpha}\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}-\frac{\partial_{q}^{2}}{\partial_{q}x_{\nu}^{2}},$ (6)
+$\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}$ $=\delta_{\mu\alpha}\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}},\quad\mu,\nu,\alpha=1,2,\dotsc,n.$ (7)
+Proof.
+The proof relies on the observation that the square of (5) is equivalent to $-\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}^{2}}-\frac{\partial_{q}^{2}}{\partial_{q}x_{\nu}^{2}}$.
+Computing $D_{q}^{2}$ gives
+$D_{q}^{2}$ $=\mathrm{e}_{\mu}^{2}\frac{\partial_{q}^{2}}{\partial_{q}x_{\nu}^{2}}+\mathrm{e}_{\nu}^{2}\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}^{2}}+\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}.$ (8)
+Substituting (6) and (7) into (8) yields
+$D_{q}^{2}$ $=-2\delta_{\mu\alpha}\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}-\frac{\partial_{q}^{2}}{\partial_{q}x_{\nu}^{2}}+\delta_{\mu\alpha}\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}.$ (9)
+For $\mu=\alpha$, we obtain the claimed identity
+$D_{q}^{2}=-\frac{\partial_{q}^{2}}{\partial_{q}x_{\mu}^{2}}-\frac{\partial_{q}^{2}}{\partial_{q}x_{\nu}^{2}},$
 which completes the proof. ∎
-Now, we will define a new $q$-differential operator for $q$-spinor variables over an orthonormal basis of $\mathds{R}^{n}$, different to usual Dirac operator and Cauchy - Riemann operator mentioned in [3], [4], [6], and [7].
-2.2 The new $q$-differential operator for $q$-spinor variables
-The motivation comes from the construction of the any diiferential operator that satisfy the property $D_{q}^{2}=-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}$ for all $\mu,\nu=1,2,...,n$ over an orthonormal basis of $\mathds{R}^{n}$.
-Proposition 2.2.
-Let $\left\{\mathrm{e}_{1},\mathrm{e}_{2},...,\mathrm{e}_{n}\right\}$ be an orthonormal basis of $\mathds{R}^{n}$. The $q$- Differential operator $D^{q}$
-$D_{q}=\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}},$ (55)
-is defined over $\mathds{R}^{n}$ under the follows relations
-$\mathrm{e}^{2}_{\mu}\frac{\partial_{q}^{2}}{\partial_{q}x^{2}_{\nu}}+\mathrm{e}^{2}_{\nu}\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}$ $=\ -2\delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}},$ (56)
-$\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}$ $=\ \delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}},\quad\mu,\nu,\alpha=1,2,...n.$ (57)
-Proof.
-The proof is based on the following observation: the square of (55) is equivalent to $-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}$.
-Therefore, we first compute $D^{2}_{q}$ resulting
-$D^{2}_{q}$ $=\ \mathrm{e}^{2}_{\mu}\frac{\partial_{q}^{2}}{\partial_{q}x^{2}_{\nu}}+\mathrm{e}^{2}_{\nu}\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}},$ (58)
-and substituting (56) and (57) into (58) we get
-$D^{2}_{q}$ $=\ -2\delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}+\delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}},$ (59)
-and for $\mu=\alpha$ finally we obtain $-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}$, which completes the proof. ∎
-According to above, we can said that this operator acts over a $q$-spinor function on the form $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))$. Therefore (55) can be expressed as
-$D_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))=\mathrm{e}_{\nu}\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\nu}}.$ (60)
-and the derivatives $\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\mu}}$ and $\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\nu}}$ can be determined using the $q$-spinor chain rule defined by (51).
-We will consider the following example
+According to the above result, this operator acts on $q$-spinor functions of the form $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))$. Therefore, the operator (5) can be rewritten as
+$D_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))=\mathrm{e}_{\nu}\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial_{q}x_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial_{q}x_{\nu}},$ (10)
+where the $q$-derivatives are computed using the chain rule (1) for $q$-spinor variables.
 Example 2.3.
-Let $\Psi(x_{\mu},x_{\nu})=\exp{ix_{\mu}}$ ) and $u^{1}_{\dot{2}}=q^{2}x_{\mu}$.
-Now, to obtain the differential operator (55), first we write $\Psi(u^{1}_{\dot{2}})$ as $\exp{iu^{1}_{\dot{2}}}$ ).
-Later we apply (51)
-$\frac{\partial_{q}\Psi}{\partial_{q}x_{\mu}}$ $=\ \frac{\partial_{q}}{\partial_{q}u^{1}_{\dot{2}}}(\exp{iu^{1}_{\dot{2}}})\frac{\partial_{q}}{\partial_{q}x_{\mu}}(q^{2}x_{\mu}),$
-[…]
-$=\ q^{2}\frac{\exp{iqu^{1}_{\dot{2}}}-q\exp{iqu^{1}_{\dot{2}}}}{[qu]^{1}_{\dot{2}}-q(u^{1}_{\dot{2}})},$
-finally we apply (55) to obtain
-$D_{q}\Psi(u^{1}_{\dot{2}})=\mathrm{e}_{\nu}q^{2}\frac{\exp{iqu^{1}_{\dot{2}}}-q\exp{iqu^{1}_{\dot{2}}}}{[qu]^{1}_{\dot{2}}-q(u^{1}_{\dot{2}})}.$
-The expression (55) can be applied to functions that not depends of $u^{\alpha}_{\dot{\beta}}$, That is, it depends only on the variables $x_{\mu},x_{\nu}$.
-To obtain this result, it is necessary to define the $q$- derivatives in terms of the $q$-spinor variable $x^{\alpha}$ and $x_{\dot{\beta}}$, and the elements $\mathrm{e}_{\mu},\mathrm{e}_{\nu}$ in the following definition
+Let $\Psi(x_{\mu},x_{\nu})=\exp(ix_{\mu})$ and define $u^{1}_{\dot{2}}=q^{2}x_{\mu}$. Then, $\Psi(u^{1}_{\dot{2}})=\exp(iu^{1}_{\dot{2}})$. Using (1), we compute:
+$\frac{\partial_{q}\Psi}{\partial_{q}x_{\mu}}$ $=\frac{\partial_{q}}{\partial_{q}u^{1}_{\dot{2}}}(\exp(iu^{1}_{\dot{2}}))\cdot\frac{\partial_{q}}{\partial_{q}x_{\mu}}(q^{2}x_{\mu})$
+$=q^{2}\cdot\frac{\exp(iqu^{1}_{\dot{2}})-q\exp(iu^{1}_{\dot{2}})}{(qu)^{1}_{\dot{2}}-qu^{1}_{\dot{2}}}\cdot\frac{\partial_{q}x_{\mu}}{\partial_{q}x_{\mu}}$
+$=q^{2}\cdot\frac{\exp(iqu^{1}_{\dot{2}})-q\exp(iqu^{1}_{\dot{2}})}{(qu)^{1}_{\dot{2}}-qu^{1}_{\dot{2}}},$
+and applying (5), we obtain:
+$D_{q}\Psi(u^{1}_{\dot{2}})=\mathrm{e}_{\nu}q^{2}\cdot\frac{\exp(iqu^{1}_{\dot{2}})-q\exp(iqu^{1}_{\dot{2}})}{(qu)^{1}_{\dot{2}}-qu^{1}_{\dot{2}}}.$
+The expression (5) also applies to functions that depend directly on the variables $x_{\mu}$ and $x_{\nu}$, without requiring the spinor composition.
+For this case, we define the $q$-derivatives in terms of the $q$-spinor variables $x^{\alpha}$ and $x_{\dot{\beta}}$ as follows:
 Definition 2.4.
-Let us consider the function $\psi(x_{\mu},x_{\nu})$. The expressions:
+Let $\psi(x_{\mu},x_{\nu})$ be a scalar function. The $q$-derivatives with respect to $x_{\mu}$ and $x_{\nu}$ are defined by
 $\frac{\partial_{q}\psi}{\partial_{q}x_{\mu}}$ $=\frac{\psi(x_{\mu}+q\,\mathrm{e}_{\mu}x^{\alpha})-\psi(x_{\mu})}{x^{\alpha}},$ (11)
 $\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}$ $=\frac{\psi(x_{\nu}+q\,\mathrm{e}_{\nu}x_{\dot{\beta}})-\psi(x_{\nu})}{x_{\dot{\beta}}},$ (12)
-are the $q$-derivatives respect to $x_{\mu}$ and $x_{\nu}$ in terms of $q$- spinor variables $x_{\dot{\beta}},x^{\alpha}$ and the elements $\mathrm{e}_{\mu},\mathrm{e}_{\nu}$.
-Let us consider the following remark.
+where $x^{\alpha}$ and $x_{\dot{\beta}}$ are $q$-spinor variables, and $\mathrm{e}_{\mu},\mathrm{e}_{\nu}$ belong to the orthonormal basis of $\mathds{R}^{n}$.
 Remark 2.5.
-Taking into account the above claim, we can said that the differential operator not apply only over the functions $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$, but also over the function $\psi(x_{\mu},x_{\nu})$, and therefore the operator (55) also depends of the variables $x_{\mu}$ and $x_{\nu}$.
+In view of the above, we conclude that the operator (5) not only acts on spinor-composed functions $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$, but also on scalar functions $\psi(x_{\mu},x_{\nu})$.
+Hence, it inherently depends on the variables $x_{\mu}$ and $x_{\nu}$.
 Example 2.6.
-Consider the following function $\psi(x_{\mu},x_{\nu})=qx_{\nu}x_{\dot{\beta}},\dot{\beta}=\dot{2}$. Now, to determine the derivatives, we apply (61) and (62)
-$\frac{\partial_{q}\psi}{\partial_{q}x_{\mu}}$ $=0$
-$\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}$ $=\ \frac{q(x_{\nu}+q\mathrm{e}_{\nu}x_{\dot{2}})x_{\dot{2}}-qx_{\nu}x_{\dot{2}}}{x_{\dot{2}}}$
-$=\ q^{2}\mathrm{e}_{\nu}x_{\dot{2}},$
-thus the operator (55) is expressed as
+Consider the scalar function $\psi(x_{\mu},x_{\nu})=qx_{\nu}x_{\dot{\beta}}$, with $\dot{\beta}=\dot{2}$. Applying (11) and (12), we obtain:
+$\frac{\partial_{q}\psi}{\partial_{q}x_{\mu}}$ $=0,$
+$\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}$ $=\frac{q(x_{\nu}+q\,\mathrm{e}_{\nu}x_{\dot{2}})x_{\dot{2}}-qx_{\nu}x_{\dot{2}}}{x_{\dot{2}}}=q^{2}\mathrm{e}_{\nu}x_{\dot{2}}.$
+Thus, the operator (5) yields:
 $D_{q}\psi=q^{2}x_{\dot{2}}\,\mathrm{e}_{\mu}\mathrm{e}_{\nu}.$
@@
 Definition 2.7.
-The $q$-analogue of (42) can be expressed as
+The $q$-analogue of the Dirac operator is defined by
 $D_{\mu}^{q}=\gamma_{\mu}\frac{\partial^{q}}{\partial^{q}x_{\mu}}.$ (13)
-We can now formulate our main results in the following propositions.
+We are now ready to state our main results in the following propositions.
 2.4 The $q$-differential operators for $q$-spinor variables
 Proposition 2.8.
-Let $\Psi$ be a function on the $q$-spinor variables. The operator (55) for $q$-spinor variables $D^{q}\Psi$ is given by
+Let $\Psi$ be a function of the $q$-spinor variables. The operator (5) for $q$-spinor variables $D^{q}\Psi$ is given by
 $D^{q}\Psi=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}u^{\alpha}_{\dot{\beta}}.$ (14)
@@
 $\frac{\partial^{q}\Psi}{\partial^{q}x}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}}.$ (15)
-Multiplying the left-hand side by $\mathrm{e}_{\nu}$ in (54), we have
+Multiplying the left-hand side by $\mathrm{e}_{\nu}$ in (4), we obtain
 $\begin{split}\mathrm{e}_{\nu}\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}&=\ \mathrm{e}_{\nu}\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}\\ &=\ \frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\mathrm{e}_{\nu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}},\end{split}$ (16)
@@
 $\begin{split}\mathrm{e}_{\mu}\frac{\partial^{q}\Psi}{\partial^{q}x_{\nu}}&=\ \mathrm{e}_{\mu}\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}}\\ &=\ \frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\mathrm{e}_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}},\end{split}$ (17)
-adding (66) with (67) and considering (55), finally results
+Adding (16) and (17) and considering (5), we finally obtain
 $D^{q}\Psi=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}u^{\alpha}_{\dot{\beta}},$
-which our assertion. ∎
+which proves our assertion. ∎
 Remark 2.9.
-The above proof implies the following relation
+The above proof implies the following relation:
 $\mathrm{e}_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}}-\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}}\mathrm{e}_{\mu}=0.$ (18)
 Remark 2.10.
-If $\Psi$ also depends of $x_{\nu}$, then
+If $\Psi$ also depends on $x_{\nu}$, then
 $D^{q}\Psi=\frac{\partial_{q}\Psi}{\partial^{q}x_{\nu}}D^{q}x_{\nu}.$ (19)
@@
 $D^{q}x_{\nu}:=\mathrm{e}_{\mu}\mathrm{d}^{q}x_{\nu},$ (20)
-consequently the $q$-differential for $u^{\alpha}_{\dot{\beta}}$ is defined as
+consequently, the $q$-differential for $u^{\alpha}_{\dot{\beta}}$ is defined as
 $D^{q}u^{\alpha}_{\dot{\beta}}:=\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\nu}}D^{q}x_{\nu}.$ (21)
 Proposition 2.12.
-Let $\Psi$ be a function on the $q$-spinor variables. The Dirac operator for $q$-spinor variables $D_{\mu}^{q}\Psi$ is given by
+Let $\Psi$ be a function of the $q$-spinor variables. The Dirac operator for $q$-spinor variables $D_{\mu}^{q}\Psi$ is given by
 $D^{q}_{\mu}\Psi=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}.$ (22)
 Proof.
-Multiplying the left-hand side by $\gamma_{\mu}$ in (54), we have
+Multiplying the left-hand side by $\gamma_{\mu}$ in (4), we obtain
 $\begin{split}\gamma_{\mu}\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}&=\ \gamma_{\mu}\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}\\ &=\ \frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\gamma_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}},\end{split}$ (23)
-and considering (63), finally we get
+and considering (13), we finally get
 $D_{\mu}^{q}\Psi=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}_{\mu}u^{\alpha}_{\dot{\beta}},$
@@
 Remark 2.13.
-The above proof implies the following relation
+The above proof implies the following relation:
 $\gamma_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}-\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}\gamma_{\mu}=0.$ (24)
@@
 $D^{q}_{\mu}x:=\gamma_{\mu}\mathrm{d}^{q}x^{\mu},$ (25)
-consequently the Dirac $q$-differential for $u^{\alpha}_{\dot{\beta}}$ is defined as
+consequently, the Dirac $q$-differential for $u^{\alpha}_{\dot{\beta}}$ is defined as
 $D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}:=\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x^{\mu}}D^{q}_{\mu}x^{\mu}.$ (26)
-Now, by (52), (64) and (72), we have the follows $q$-spinor integral formulas showed in the following subsection.
 2.5 The integral formulas in $q$-spinor variables
@@
 $\oint_{\Gamma_{q}}\frac{\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{qu^{\alpha}_{\dot{\beta}}(x_{\mu})-(qu)^{\alpha}_{\dot{\beta}}(x_{0})}$ $=\ \frac{1}{q}\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n},$ (28)
-where $\gamma_{\mu}$ are the $q$-deformed Dirac matrices defined in (43).
+where $\gamma_{\mu}$ are the $q$-deformed Dirac matrices defined in the reference [14].
 Proof.
-The proof of this result is based on the ideas presented in [8]. Combining (38) with (72) we get
+The proof of this result follows from the approach developed in [8], consequently, expression (22) takes the form
 $D^{q}_{\mu}\Psi=\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}-qu^{\alpha}_{\dot{\beta}}}-\frac{q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}-qu^{\alpha}_{\dot{\beta}}},$ (29)
@@
 $\oint_{\Gamma_{q}}D^{q}_{\mu}\Psi=\oint_{\Gamma_{q}}\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})}-\oint_{\Gamma_{q}}\frac{q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})}.$ (30)
-Hence, to solve the integral $\oint_{\Gamma_{q}}D^{q}_{\mu}\Psi$, we will use similary the proof of the Theorem 2.9 of the reference [8], interchanging the Pauli matrices of $q$-deformed Minkowski space (e.g.
-[14]) by the $q$-deformed Dirac matrices (43), obtaining
+Hence, to solve the integral $\oint_{\Gamma_{q}}D^{q}_{\mu}\Psi$, we will use similary the proof of the Theorem 2.9 of the reference [8], interchanging the Pauli matrices of $q$-deformed Minkowski space by the $q$-deformed Dirac matrices (e.g.
+[14]) , obtaining
 $\begin{split}\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}-\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}&=\\ \oint_{\Gamma_{q}}\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})}-\oint_{\Gamma_{q}}\frac{q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})},\end{split}$ (31)
 and finally, equalating terms that depend on $\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu})$ and $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$ we obtain (27) and (28), and the proof is complete. ∎
-We will mention an important consequence of the Definition 3.4.
-starting of Eqs.
+We will mention an important consequence of the Proposition 2.15 starting of Eqs.
 (27) and (28): the formulation of the integral $\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}}(x))d^{q}x_{\mu}$, which we will mention in the following theorem.
@@
 $\frac{\partial^{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D_{\mu}^{q}u^{\alpha}_{\dot{\beta}}=b\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu},$ (35)
-from (38) it follows that
+the term $\frac{\partial^{q}\Psi(u^{\alpha}{\dot{\beta}})}{\partial^{q}u^{\alpha}{\dot{\beta}}}$ may be written in the following form
 $\frac{\partial^{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial^{q}u^{\alpha}_{\dot{\beta}}}=\dfrac{q\Psi(u^{\alpha}_{\dot{\beta}})}{qu^{\alpha}_{\dot{\beta}}-(qu)^{\alpha}_{\dot{\beta}}},$ (36)
@@
 $D^{q}_{\mu}\psi(u^{\alpha}_{\dot{\beta}})-b\phi(u^{\alpha}_{\dot{\beta}})$ = 0,absent 0=\ 0,= 0 , (42)
-$D^{q}\psi(u^{\alpha}_{\dot{\beta}})-a\phi(u^{\alpha}_{\dot{\beta}})$ = 0absent 0=\ 0= 0 (93)
+$D^{q}\psi(u^{\alpha}_{\dot{\beta}})-a\phi(u^{\alpha}_{\dot{\beta}})$ = 0.absent 0=\ 0.= 0 . (43)
 which we will show in the following section.
@@
 Remark 3.2.
-We first consider the new $q$-differential operator given by (55).
-Now, to solve the integral $\oint_{\Gamma_{q}}D^{q}\Psi$, we will use similary the proof of the Theorem 2.9 of the reference [8] (also can see the proof of Proposition 3.10.) to obtain similar expressions to (96) and (97), i.e.
+We begin by considering the new $q$-differential operator defined in (5).
+In order to evaluate the integral $\oint_{\Gamma_{q}}D^{q}\Psi$, we shall proceed in a manner analogous to the proof of Theorem 2.9 in [8] (see also the proof of Proposition 3.10 therein).
+This approach allows us to derive expressions similar to (46) and (47), namely:
 $\oint_{\Gamma_{q}}D^{q}\psi(u^{\alpha}_{\dot{\beta}}(x_{0}))$ $=\psi(u^{\alpha}_{\dot{\beta}}(x_{0})),$ (48)
@@
 To obtain $D^{2}_{q}$ it is necessary to use the relations (6) and (7).
-This operator is different to usual Dirac operator and Cauchy Riemann operator mentioned in [3], [4], [6], and [7].
+This operator differs from the classical Dirac and Cauchy–Riemann operators discussed in [3, 4, 6, 7, 9], and may be regarded as a $q$-deformed version of the operator $D$ introduced in [9].
 In the case of the Dirac operator for $q$-spinor variables, defined by (22), has been stablished from the Remarks 2.10 and 2.11.
@@
 $\int_{\Omega_{q}}\frac{\psi[(1-q^{-1})\textit{{p}}_{\dot{k}}]\mathbf{d}_{q}\textit{{p}}_{\dot{k}}}{\textit{{p}}_{\dot{k}}-\textit{{u}}^{\beta}_{\dot{\alpha}}}$ $=\ \sum\limits_{n=0}^{\infty}[\gamma^{\mu}\gamma^{\nu}\psi[(1-q^{-1})\textit{{u}}^{\beta}_{\dot{\alpha}}]]^{n}.$ (78)
+The formulation of Dirac operators in $q$-deformed spinorial coordinates constitutes a significant development in the study of quantum theories defined on noncommutative spaces.
+These structures enable a rigorous extension of the foundations of quantum mechanics and quantum field theory to contexts where classical continuous symmetries are replaced by quantum symmetries, described in terms of Hopf algebras and quantum groups.
+Even in its classical form, the Dirac operator plays a central role in the description of spin-$\frac{1}{2}$ particles, such as electrons, and provides a relativistic formulation of quantum mechanics.
+Its generalisation to the $q$-deformed framework incorporates effects arising from discretisation, quantum curvature, and underlying noncommutative geometries.
+This generalisation is particularly relevant in theoretical approaches that seek to unify quantum mechanics with gravitational phenomena.
+From an algebraic perspective, $q$-deformed Dirac operators are intimately connected with representations of $q$-deformed Clifford algebras and with noncommutative geometries in the sense of Connes.
+Their study permits the characterisation of quantum manifolds equipped with spinorial structure, thereby paving the way for more general formulations of field theory on curved or quantum spacetimes.
+In physical terms, these operators facilitate the modelling of fermionic dynamics in the presence of $q$-deformed external fields, including noncommutative electromagnetic potentials and generalised gauge fields.
+Notably, the inclusion of a $q$-deformed Maxwell algebra introduces novel forms of interaction between fields and particles, with potential applications to effective models of quantum gravity, quantum cosmology, and condensed matter systems exhibiting discrete or topological symmetries.
+The most substantial contribution lies in the capacity of these operators to encapsulate, within a single mathematical framework, the interplay of quantum symmetries, generalised spinorial structures, and noncommutative dynamics.
+As such, they offer a powerful and versatile tool for probing novel physical regimes in which the geometry of spacetime itself is governed by quantum principles.